### v1

# Benchmarking Mental State Representations in Language Models

Abstract
While numerous works have assessed the generative performance of language models (LMs) on tasks requiring Theory of Mind reasoning, research into the models’ internal representation of mental states remains limited. Recent work has used probing to demonstrate that LMs can represent beliefs of themselves and others. However, these claims are accompanied by limited evaluation, making it difficult to assess how mental state representations are affected by model design and training choices. We report an extensive benchmark with various LM types with different model sizes, fine-tuning approaches, and prompt designs to study the robustness of mental state representations and memorisation issues within the probes. Our results show that the quality of models’ internal representations of the beliefs of others increases with model size and, more crucially, with fine-tuning. We are the first to study how prompt variations impact probing performance on theory of mind tasks. We demonstrate that models’ representations are sensitive to prompt variations, even when such variations should be beneficial. Finally, we complement previous activation editing experiments on Theory of Mind tasks and show that it is possible to improve models’ reasoning performance by steering their activations without the need to train any probe.
Machine Learning, ICML
1 Introduction
Modern language models (LMs) trained on next token prediction have demonstrated impressive capabilities, spanning coding, mathematical reasoning, fact verification, and embodied interaction (Wei et al., 2022; Bubeck et al., 2023). As these models are designed with the ultimate goal of collaborating with humans, it becomes imperative that they complement these skills with an understanding of humans, in particular their beliefs, emotions, desires, and intentions (Li et al., 2023a). Core to this understanding is Theory of Mind (ToM) – the ability to attribute mental states to oneself and others (Premack & Woodruff, 1978). ToM is essential for effective communication and cooperation with other agents, facilitating interaction and learning from feedback and demonstrations (Saha et al., 2023). Given its significance, ToM has emerged as a critical milestone in artificial intelligence (AI) and an important capability when evaluating cutting-edge LMs (Bubeck et al., 2023). Interest in LMs’ generative performance on tasks requiring ToM reasoning has resulted in a wide variety of benchmark datasets, typically involving question-answering tasks (Le et al., 2019; Gandhi et al., 2023; Kim et al., 2023; He et al., 2023; Tan et al., 2024; Xu et al., 2024).
Despite showing improved performance on ToM benchmarks compared to earlier models, modern LMs are still far from perfect (Sap et al., 2022). Text generated by LMs often contains errors that limit their performance on ToM tasks (Martindale et al., 2019). Previous work has shown that it is sometimes possible to still obtain correct predictions by probing LMs’ internal representations (Li et al., 2021; Liu et al., 2023b; Gurnee et al., 2023). In particular, Zhu et al. (2024) have shown that LMs, when prompted with a story and a belief statement, can represent beliefs from their own perspective and, to a lesser extent, from the perspective of a character in the story. Their work is an important first step towards understanding how LMs represent mental states, but it is limited in the number of models and settings studied, leaving many questions unanswered.
Building and extending on (Zhu et al., 2024), we benchmark mental state representations of self and others in language models through extensive experiments of different LM families, model sizes, and fine-tuning approaches. Specifically, we design a set of experiments to address the following research questions: RQ1. What is the relation between model size and probing accuracy? RQ2. Does fine-tuning with instruction-tuning (Wei et al., 2021) and/or reinforcement learning from human feedback (Christiano et al., 2017; Ouyang et al., 2022, RLHF) have an effect on probing accuracy? RQ3. Are models’ internal representations of beliefs sensitive to prompt variations? RQ4. Is there a risk of probes memorising training data due to the large dimensionality of LM representations? RQ5. Can we enhance LMs’ performance by editing their activations without training dedicated probes?
To answer RQ1, we perform probing experiments on two families of LMs, Llama-2 (Touvron et al., 2023), and Pythia (Biderman et al., 2023), ranging from models with 70 million to 70 billion parameters. To address RQ2, we compare the probing performance of models pre-trained solely on next token prediction with models that have been fine-tuned using instruction-tuning and/or RLHF. Our experiments reveal that probing accuracy on the beliefs of others increases with model size and, more crucially, with fine-tuning. To answer RQ3, we repeat our probing experiments using different variations of the prompt used by Zhu et al.. Specifically, we consider two variations that are expected to negatively impact LMs’ representations (Random, Misleading), and two that are supposed to have a positive influence (Time Specification, Initial Belief). By conducting these experiments, our work is the first to explore the sensitivity of LMs’ representations to prompting in the context of ToM. Our findings demonstrate that models’ representations are sensitive to prompt variations, even when such variations should be beneficial. To address RQ4, we compare our trained probes with a second set of probes trained only on the representations’ first top $k$ principal components. This requires learning much fewer parameters and eliminates any possible memorisation issue. We find no strong evidence of memorisation in the probes, as it is possible to recover most of the accuracy by training probes on a small subset of principal components of models’ representations. We formulate RQ5 as a follow-up question to Zhu et al. (2024) who found that probes trained to predict beliefs can be used to steer models’ activation using inference-time intervention (Li et al., 2023c, ITI) to improve LMs’ downstream performance on ToM tasks. In contrast, we show that by using contrastive activation addition (Rimsky et al., 2023, CAA), we can steer models’ activations without the need to train any probe and, in a more generalisable way, obtain significant performance improvements across different ToM tasks.
In summary, our work makes the following contributions:
  1. 1.
    We report extensive probing experiments with various types of LMs with different model sizes and fine-tuning approaches, showing that the quality of models’ internal representations of the beliefs of others increases with model size and, more crucially, fine-tuning.
  2. 2.
    We are the first to study how prompt variations impact belief probing performance, showing that models’ representations are sensitive to prompt variations, even when such variations should be beneficial.
  3. 3.
    We show that by using contrastive activation addition it is possible to improve models’ reasoning performance by steering their activations without the need to train any probe.
2 Related Work
Machine Theory of Mind
Theory of Mind (ToM) has been studied in cognitive science and psychology for decades (Gurney et al., 2021). Mirroring efforts to understand ToM in humans, an increasing number of works in the computational sciences have investigated means to equip AI with similar capabilities. Previously proposed models that aim to implement a machine ToM have been based on partially observable Markov decision processes (POMDP) (Doshi et al., 2010; Han & Gmytrasiewicz, 2018), Bayesian methods (Baker et al., 2011, 2017) and deep learning methods (Rabinowitz et al., 2018; Bara et al., 2021; Wang et al., 2022; Duan et al., 2022; Liu et al., 2023a; Bortoletto et al., 2024b, a). Recent advances in LMs have sparked interest in evaluating their ToM capabilities. Various benchmarks have been proposed, aiming to measure LMs’ ability to understand and reason about the beliefs, goals, and intentions of others (Le et al., 2019; He et al., 2023; Kim et al., 2023; Gandhi et al., 2023; Xu et al., 2024; Tan et al., 2024). Additionally, efforts have been made to enhance LMs’ ToM through prompting techniques (Zhou et al., 2023b; Moghaddam & Honey, 2023; Wilf et al., 2023). A new direction of research explores LMs’ internal representation of mental states. Zhu et al. (2024) demonstrated that LMs linearly encode beliefs from different agents’ perspectives, and manipulating these representations can enhance ToM task performance. While Zhu et al.’s work is a crucial initial step, our work dives deeper into LMs’ internal belief representations, offering a broader insight into these mechanisms.
Probing Neural Representations
Initially proposed by Alain & Bengio (2017), probing has emerged as a common method for determining if models represent particular features or concepts. In the realm of LMs, numerous works used probing to demonstrate that these models acquire rich linguistic representations. These representations span syntactic and semantic concepts such as syntactic categories, dependency relations, co-reference, and word meaning (Conneau et al., 2018; Tenney et al., 2018, 2019; Rogers et al., 2021; Li et al., 2021; Hernandez & Andreas, 2021; Marks & Tegmark, 2023; Liu et al., 2023b). A separate line of work explored if and how LMs represent the world, i.e., whether they possess a world model. Li et al. (2021) showed that LMs track the states of entities within a context. Other works showed that LMs exhibit representations reflecting non-linguistic concepts in the world, which LMs have never observed (Abdou et al., 2021; Patel & Pavlick, 2022; Li et al., 2023b; Nanda et al., 2023). An emergent line of work that is particularly relevant to our work used probing to explore if LMs have agent models, for example, if they can represent beliefs of self and others (Zhu et al., 2024; Bortoletto et al., 2024a). While representing an important first step towards understanding the internals of Theory of Mind in LMs, experiments in (Zhu et al., 2024) are limited in settings and models considered. In this work, we contribute with extensive experiments that employ a wider variety of LMs and a wider range of settings. Furthermore, we also explore possible memorisation issues in the probes.
Prompt Analysis
Research on prompt robustness in LMs is still in its infancy but has quickly sparked much interest. On one hand, previous work has shown that LMs are vulnerable to prompt alterations like token deletion or reordering (Ishibashi et al., 2023), biased or toxic prompts (Shaikh et al., 2023) and similarity to training data (Razeghi et al., 2022). On the other hand, instruction-tuned models have proved to be more robust against prompt variation, even when using misleading instructions (Webson & Pavlick, 2022). Other works have shown the importance of input-output format (Min et al., 2022) and of demonstration example ordering for few-shot performance (Zhao et al., 2021; Lu et al., 2022; Zhou et al., 2023a). In this work, we shift our focus from analysing how sensitive model outputs are to how model representations change. Our work, along with (Gurnee et al., 2023), is one of the first to explore how prompt design affects how accurately models represent concepts. In particular, Gurnee et al. (2023) have studied whether LMs’ representations of space and time are robust to prompt variations. In stark contrast, we explore for the first time the effect of prompt variations on how models represent mental states internally.
Activation Editing
Recent advancements in NLP have introduced innovative techniques for controlling and manipulating text generation models. While weight editing proposed to modify models’ weights (Meng et al., 2022; Ilharco et al., 2022; Orgad et al., 2023), activation editing has emerged as an alternative way to influence model behaviour without any additional fine-tuning (Li et al., 2023b; Hernandez et al., 2023). This approach involves manipulating the internal representations of models to direct their outputs towards desired outcomes. One notable method in this domain is inference-time intervention (Li et al., 2023c, ITI), which has been proposed to enhance truthfulness in LMs. ITI involves training linear probes on contrastive question-answering datasets to identify “truthful” attention heads and then shifting attention head activations during inference along the identified truthful directions. In contrast, activation addition (Turner et al., 2023, AA) and contrastive activation addition (Rimsky et al., 2023, CAA) offer ways to generate steering vectors by only using LMs’ intermediate activations. Zhu et al. have used ITI to show that it is possible to manipulate LMs’ internal representations of mental states (Zhu et al., 2024). In this work, we show that using CAA can further improve LMs’ ToM capabilities without the necessity of training any probe. Remarkably, CAA operates at the residual stream level, eliminating the need for a fine-grained search over attention heads.
3 Experimental Setup
3.1 Probing
Figure 1: Example of false belief from our probing datasets. The labels $z_{p}$ and $z_{o}$ correspond to $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$, respectively. By manipulating the protagonist’s percepts after the causal event we obtain two scenarios: true belief and false belief.
In line with previous work (Zhu et al., 2024) we linearly decode belief status from the perspective of different agents by using probing (Alain & Bengio, 2017). Probing involves localising specific concepts in a neural model by training a simple classifier (called a probe) on model activations to predict a target label associated with the input data. To provide a formal definition, we adopt a similar notation to the one introduced in (Belinkov, 2022). Let us define an original model $f:x\mapsto\hat{y}$ that is trained on a dataset $\mathcal{D}^{O}=\{x^{(i)},y^{(i)}\}$ to map input $x$ to output $\hat{y}$. Model performance is evaluated by some measure, denoted $\textsc{Perf}(f,\mathcal{D}^{O})$. A probe $g:f_{l}(x)\mapsto\hat{z}$ maps intermediate representations of $x$ in $f$ at layer $l$ to some property $\hat{z}$, which is the label of interest. The probe $g$ is trained on a probing dataset $\mathcal{D}^{P}=\{x^{(i)},z^{(i)}\}$ and evaluated using some performance measure $\textsc{Perf}(g,f,\mathcal{D}^{O},\mathcal{D}^{P})$. In our case, $f$ is an autoregressive language model that given a sequence of tokens $x$ outputs a probability distribution over the token vocabulary to predict the next token in the sequence. Our probe is a logistic regression model $\hat{z}=\sigma(Wa_{l}+b)$ trained on neural activations $f_{l}(x)=a_{l}$ to predict belief labels $y=\{0,1\}$.
3.2 Dataset
Following Zhu et al. (2024) we use the BigToM benchmark (Gandhi et al., 2023). BigToM is constructed using GPT-4 (Achiam et al., 2023) to populate causal templates and combine elements from these templates. Each causal template is set up with a context and a description of the protagonist (e.g. “Noor is working as a barista […]”), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”). The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”). The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs. In this work, we focus on the Forward Belief setting proposed by (Zhu et al., 2024) in which models have to infer the belief of the protagonist given the percepts of the causal event, $P(\mathrm{belief}|\mathrm{percepts})$. We report additional details in Section A.1.1
Probing datasets
We consider two probing datasets: $\mathcal{D}_{p}^{P}=\{x^{(i)}_{p},z^{(i)}_{p}\}$, where the labels $z^{(i)}_{p}$ correspond to ground-truth beliefs from the protagonist perspective, and $\mathcal{D}_{o}^{P}=\{x^{(i)}_{o},z^{(i)}_{o}\}$, where the labels $z^{(i)}_{o}$ reflect the perspective of an omniscient oracle. $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ are built by pairing each story in BigToM with a belief statement, as shown in Figure 1. After prompting the model with a story-belief pair we cache the residual stream activations at the final token position for all residual streams (Figure 5).
3.3 Models
Zhu et al. (2024) have used two models for their experiments: Mistral-7B-Instruct (Jiang et al., 2023) and DeepSeek-7B-Chat (Bi et al., 2024) – both being the same size and fine-tuned. In contrast, we study two families of LMs that offer us options in model sizes and fine-tuning: Pythia (Biderman et al., 2023) and Llama-2 (Touvron et al., 2023). While Llama-2 offers “chat” versions fine-tuned using supervised learning and RLHF, Pythia’s open-source training set (Gao et al., 2020) ensures that there is no data leakage [^1] Llama-2 was released later than BigToM.. Additionally, we consider a version of Pythia-6.9B fine-tuned on a mixture of open-source instruction datasets (Wang et al., 2024), which we refer to as Pythia-6.9B-chat. [^2] https://huggingface.co/allenai/open-instruct-pythia-6.9b-tulu A summary of the models we study is reported in Table 2.
3.4 Probing Experiments
We aim to contribute to understanding how LMs represent beliefs of self and others by proposing a set of extensive probing experiments across LMs that differ in architecture, size, and fine-tuning approach. Our approach is generally similar to the one used by (Zhu et al., 2024), but we make a different operational choice: While (Zhu et al., 2024) train probes on each attention head for every layer, we train probes on the residual stream for every layer. We opted to use the residual stream as it integrates information from both the attention and feed-forward components, potentially encoding richer representations. Additionally, since the residual activations directly contribute to the final output predictions, probing them may better align with understanding the model’s behaviour for downstream tasks.
Model Size and Fine-tuning
We first report experiments to better understand the effect of model size and fine-tuning on belief probing accuracy. Specifically, we ask the following questions: Is there a relation between model size and probing accuracy? (RQ1) Does fine-tuning an LM with instruction-tuning or RLHF have an effect on probing accuracy? (RQ2) To answer these questions we performed the same probing experiment across all our models and compared the results.
Figure 2: Belief probing accuracy across models with different architecture, size and fine-tuning.
Figure 3: Sensitivity of protagonist belief probing accuracy to different prompt variations.
Figure 4: To investigate potential memorisation in the probes, we compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components. For Llama2: All(7b) = $4096$, All(13b) = $5120$, All(70b) = $8192$. For Pythia: All(70m) = $512$, All(410m) = $1024$, All(1b) = $2048$, All(6.9b) = $4096$, All(12b) = $5120$. We report results for Protagonist beliefs. Results for Oracle are shown in Figure 8.
Table 1: Comparison of the effects of ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023) activation editing methods on three tasks from BigToM (Gandhi et al., 2023). TB denotes a true belief task, whereas FB denotes a false belief task. The numbers represent accuracy scores, with the difference in performance compared to no intervention (No int.) indicated as subscripts (ITI -- No int. and CAA -- No int.). An asterisk (*) denotes a statistically significant difference from No int. based on a McNemar’s test (McNemar, 1947) with $p<0.05$.
Model Method Forward Belief Forward Action Backward Belief
TB FB TB $\land$ FB TB FB TB $\land$ FB TB FB TB $\land$ FB
Llama-2-7b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $44_{+0}$ $44_{+0}$ $44_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
CAA $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $74_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+30}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
Llama-2-7b-chat No int. $56$ $56$ $55$ $69$ $55$ $37$ $56$ $56$ $55$
ITI $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $69_{+0}$ $55_{+0}$ $37_{+0}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$
CAA $70_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}$ $72_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $69_{+0}$ $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $53_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $84_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}^{*}$
Llama-2-13b No int. $52$ $44$ $35$ $59$ $50$ $37$ $46$ $49$ $33$
ITI $52_{+0}$ $45_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $35_{+0}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$ $61_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $46_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $42_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$
CAA $85_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+33}}^{*}$ $88_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+44}}^{*}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+31}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $69_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}^{*}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+29}}^{*}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+43}}^{*}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+26}}^{*}$
Llama-2-13b-chat No int. $84$ $56$ $47$ $78$ $51$ $38$ $72$ $48$ $31$
ITI $84_{+0}$ $65_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $78_{+0}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $47_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}^{*}$ $72_{+0}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}$
CAA $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+38}}^{*}$ $91_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+44}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+20}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+46}}^{*}$ $87_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+56}}^{*}$
Llama-2-70b No int. $90$ $87$ $78$ $93$ $52$ $48$ $73$ $53$ $32$
ITI $90_{+0}$ $90_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $78_{+0}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $50_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $77_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+4}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$ $37_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$
CAA $99_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $95_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}^{*}$ $73_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+21}}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+39}}^{*}$ $83_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+51}}^{*}$
Llama-2-70b-chat No int. $69$ $75$ $56$ $86$ $56$ $52$ $63$ $59$ $52$
ITI $69_{+0}$ $76_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $86_{+0}$ $56_{+0}$ $52_{+0}$ $63_{+0}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$
CAA $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+23}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $89_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+32}}^{*}$ $87_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}^{*}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+8}}^{*}$ $88_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+33}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}$
Pythia-70m No int. $41$ $41$ $37$ $46$ $45$ $41$ $44$ $41$ $37$
ITI $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+8}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}$
CAA $62_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+21}}^{*}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$
Pythia-410m No int. $48$ $45$ $45$ $44$ $44$ $44$ $44$ $47$ $44$
ITI $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $62_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $52_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$
CAA $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $61_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $69_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$
Pythia-1b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
CAA $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $62_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
Pythia-6.9b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $45_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $44_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
CAA $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$
Pythia-6.9b-chat No int. $55$ $54$ $28$ $36$ $64$ $20$ $44$ $67$ $30$
ITI $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $54_{+0}$ $28_{+0}$ $44_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+8}}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $32_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $44_{+0}$ $67_{+0}$ $30_{+0}^{*}$
CAA $68_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $65_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+29}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}^{*}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}^{*}$ $67_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+24}}^{*}$
Pythia-12b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
CAA $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+20}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+23}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
Sensitivity to Prompting
By using a single prompt design, previous work left the impact of prompt design on probing accuracy unclear (Zhu et al., 2024). Our second set of experiments aims to explore how belief representations are sensitive to different prompts. Research on prompt robustness in language models is still in its infancy and focused mainly on revealing vulnerability to prompt alternations on downstream performance (Min et al., 2022; Ishibashi et al., 2023; Shaikh et al., 2023; Leidinger et al., 2023). In contrast, we study how the input influences models’ representations by asking: Are models’ internal belief representations robust to prompt variations? (RQ3) To answer this question we define four prompt variations:
  • Random: Following (Gurnee & Tegmark, 2024), we add 10 random tokens to the belief statement.
  • Misleading: Each story is followed by two belief statements, one pertinent to the story and one randomly chosen from another.
  • Time Specification: The prompt specifies that the belief statement refers to the end of the story. We study this variation because some belief statements can be true (false) at the story’s beginning but false (true) at the end. For example, consider the story in Figure 1: if Noor does not witness the swap, in the end, she will believe the pitcher contains almond milk ($y_{p}=\mathrm{True}$). However, if the same belief is referred to at the beginning of the story, then it is false ($y_{p}=\mathrm{False}$).
  • Initial Belief: We explicitly reveal the protagonist’s initial belief (e.g. “Noor believes that the pitcher contains oat milk”) in the story to test whether it biases the representations of LMs.
While all maintaining conceptual and semantic parity with the Original prompt used in (Zhu et al., 2024), Random and Misleading are expected to negatively impact LMs’ representations, while Time Specification and Initial Belief are supposed to have a positive influence. Robust representations of mental states should exhibit minimal sensitivity to these alterations. Our experiments compare probe accuracy across different model sizes, fine-tuning, and prompt variations. Examples of prompts are reported in Section A.1.4.
Memorisation
Although linear, our probes possess many learnable parameters – up to $16,385$ for Llama-2-70B. In principle, this allows them to engage in significant memorisation (Alain & Bengio, 2017). Our final set of probing experiments answers the following question: Are the probes memorising their training data? (RQ4) To answer this question, before training the probes, we project the probing datasets $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ onto their $k$ largest principal components using PCA to obtain probes with substantially fewer parameters.
3.5 Contrastive Activation Addition
Our final set of experiments builds upon the findings of (Zhu et al., 2024), who showed that employing trained probes with inference time intervention (Li et al., 2023c, ITI) could enhance LMs’ performance on ToM tasks. We take a step further and ask: Can we enhance LMs’ performance by manipulating their activations without the need for training dedicated probes? (RQ5) To find an answer we use contrastive activation addition (Rimsky et al., 2023, CAA), an extension of activation addition (Turner et al., 2023, AA) that computes steering vectors to control LMs’ behaviour. Steering vectors are computed as the average difference in residual stream activations between pairs of positive and negative instances of a specific behaviour. Formally, given a dataset $\mathcal{D}$ of triplets ($p$, $c_{p}$, $c_{n}$), where $p$ is a prompt, $c_{p}$ is a positive completion, and $c_{n}$ is a negative completion, CAA computes a mean difference vector $v^{md}_{l}$ for layer $l$ as:
$v^{md}_{l}=\frac{1}{\absolutevalue{\mathcal{D}}}\sum_{p,c_{p},c_{n}}a_{l}(p,c_{p})-a_{l}(p,c_{n})$
During inference, these steering vectors are multiplied with an appropriate coefficient $\alpha$ and added at every token position of the generated text after the user’s prompt. CAA has two main advantages over ITI: First, it eliminates the need to train probes. Second, it operates at the residual stream level, making it easier to use than methods that intervene on specific attention heads like ITI. While CAA has been used to control alignment-relevant behaviour, such as hallucinations, refusal, and sycophancy (Rimsky et al., 2023), we are the first to apply it to enhance LMs’ ToM reasoning. This can be understood as isolating the direction in the LMs’ latent space corresponding to taking the perspective of another agent. To evaluate both base and fine-tuned LMs, we rank their answers to the ToM questions according to $p_{LM}(a|q)$ (Petroni et al., 2019). We adopt the Forward Belief task split used in (Zhu et al., 2024) to compute the steering vectors. Additionally, we evaluate the transferability of the CAA steering vectors by applying them to two other BigToM tasks: Forward Action and Backward Belief. We provide details about these tasks in Section A.1.1, and a more detailed explanation of how ITI works in Section A.5.
4 Results
4.1 Effect of Model Size and Fine-tuning
Results from our study on model size and fine-tuning are shown in Figure 2. When considering oracle beliefs, probing accuracy rapidly converges to $100$, with larger models showing faster convergence rates. The smallest Pythia-70m that performs slightly worse but still achieves $95\%$ accuracy despite having less than $0.6\%$ of the parameters of Pythia-12B. This finding suggests that even small LMs can effectively represent beliefs from an omniscient perspective.
For protagonist beliefs, accuracy also increases with model size, although there is a performance gap between Llama-2 and Pythia. For example, Llama2-13B reaches around $80\%$, while Pythia-12B achieves approximately $60\%$. This gap is likely due to Llama-2 being trained on nearly seven times more tokens than Pythia. The figure also shows that accuracy at early layers is particularly low across all models. We speculate that this is due to the initial coding strategy of LMs that uses the first layers to combine individual tokens into more semantically meaningful representations (Gurnee et al., 2023). Probes on fine-tuned LMs show significantly better accuracy with improvements of up to $29\%$ for Llama2-7B-chat and $26\%$ for Pythia-6.9B-chat with respect to their base version. Fine-tuned 7B LMs outperform (Llama-2) or are on par (Pythia) with twice as large base models (12/13B), highlighting the importance of fine-tuning in developing representations of others’ beliefs. This resonates with cognitive psychology findings that ToM development is closely linked to social communication (Tomasello, 2010; Sidera et al., 2018; Ma et al., 2023), which instruction-tuning and RLHF may help induce in language models. For larger LMs, the improvements from fine-tuning decrease as model size increases (Figure 6(a)). We characterise the relationship between probe accuracy and model size in Figure 6, where we consider the best probe accuracy for every LM, i.e. the highest accuracy among probes $\{g_{l}\}$ trained on $\{a_{l}\}$ for a LM $f$. For Llama-2 base, the best probe accuracy scales logarithmically with model size ($R^{2}=0.98$, Figure 6(b)), whereas for fine-tuned it scales linearly ($R=1.0$, cf. Figure 6(c)). For Pythia base, the best probe accuracy also scales logarithmically with model size ($R^{2}=0.96$, Figure 6(d)).
4.2 Sensitivity to prompting
Figure 3 compares protagonist probe accuracy across various prompt variations for different models, considering their architecture, size, and fine-tuning. As can be seen from the figure, providing the protagonist’s Initial Belief in the story yields higher probe accuracy compared to the Original prompt (Figure 1). Accuracy for all the other prompt variations is generally lower than Original. On one hand, misleading prompts hurt performance across all models. This finding resonates with Webson & Pavlick (2022) who found that instruction-tuned models, despite being more robust, are still sensitive to misleading prompts. On the other hand, Time Specification unexpectedly does not help in disambiguating belief states in different time frames, as we hypothesised in §3.4. Additionally, models show sensitivity to Random tokens placed before the belief statement. Results for oracle beliefs are reported in Figure 7 and indicate that models maintain high accuracy. Misleading prompts slightly reduce performance to around $95\%$. In summary, these experiments show that LMs possess robust belief representations when taking an omniscient perspective, whereas their representations of others’ beliefs are more susceptible to prompt variations.
4.3 Memorisation Effects in the Probes
Figure 4 and Figure 8 show probe accuracies obtained by training a probe on the top $k$ principal components of the intermediate representations for protagonist and oracle, respectively. Specifically, we consider $k=\{2,10,100,1000\}$, spanning several orders of magnitude. For models with hidden dimensions smaller than $1000$, we skip this value. For all models, it is generally possible to recover most of the original accuracy by training probes on a number $k$ of principal components of the activations that is more than one order of magnitude smaller, indicating no strong evidence of memorisation in the probes.
4.4 Contrastive Activation Addition
We finally compare models’ accuracy on three BigToM tasks in Table 1. Each model has been evaluated three times: without any intervention, using ITI, and using CAA. Hyperparameter details can be found in Section A.6. Note that we use steering vectors computed using the Forward Belief task for all three tasks to test their generalisability (Zhu et al., 2024). As can be seen from the table, performance without intervention is generally lower across tasks and model sizes, with the larger Llama-2-70B and Llama-2-70B-chat models exhibiting higher accuracy. Performance for Pythia models of different sizes does not change much, with the fine-tuned Pythia-6.9B-chat showing better performance on single true belief (TB) and false belief (FB) tasks but not on their conjunction (Both). ITI demonstrates modest improvements over no intervention for Llama-2 models. Improvements for Pythia models are consistent and higher, up to $+17$. The only exception is Pythia-6.9B-chat, for which ITI is not always beneficial.
CAA consistently delivers the most substantial accuracy improvements across all models and tasks, up to $+56$ for Llama-2-13B-chat on the (Backward Belief), which Gandhi et al. have identified as the hardest task. Despite its relatively small size, Llama-2-13B-chat excels in all three tasks when using CAA. Larger 70B models often achieve accuracies close to or exceeding $90\%$. Smaller models like Pythia-70M and Pythia-410M also show significant gains with CAA, though the absolute performance is still lower than Llama-2. Overall, our results indicate that it is possible to effectively enhance ToM reasoning in LMs without needing to train any probe, which yields even improved results. Furthermore, we show that CAA steering vectors generalise well, yielding substantial performance gains across all ToM tasks.
5 Limitations and Future Work
Our study focused on expanding experiments from the model perspective, examining architectures, sizes, fine-tuning, and prompt design, all within the same dataset. A natural extension of our work is replicating these experiments across multiple datasets and more model families. Given the rapid pace of new language model releases, studying all available models is impractical, particularly considering computational resource constraints. Nevertheless, our approach can be adopted to support new benchmarks or to evaluate newly released models as they become available.
6 Conclusion
Our study addresses a significant gap in understanding LMs by investigating their internal representation of mental states. We conducted an extensive benchmark involving various LM types, sizes, fine-tuning approaches, and prompt designs to examine the robustness of these representations. Our findings reveal that scaling LMs’ size and, in particular for smaller LMs, fine-tuning are key to developing representations of others’ beliefs. We are the first to demonstrate that such prompt variations influence model representations, and we also demonstrate the feasibility of enhancing models’ ToM reasoning by steering their activations without training any probe. Overall, our work contributes valuable insights into the factors influencing LMs’ mental state representations, shedding light on avenues for improving their performance in ToM tasks.
Societal Impact
While our work is foundational and remains distant from specific applications with direct societal impact, it’s important to recognise the ethical implications of modelling and predicting mental states. Handling sensitive aspects of individuals’ inner experiences and emotions requires careful consideration to avoid reinforcing biases or misunderstanding psychological nuances.
Author Contributions Statement
M. Bortoletto designed and directed the project, performed the experiments, analysed the results, and wrote the manuscript.
C. Ruhdorfer provided feedback on the experimental results and contributed to editing the final manuscript.
L. Shi contributed to proofreading the final manuscript.
A. Bulling supervised the project and contributed to the writing of the manuscript.
Acknowledgements
L. Shi was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer. The authors would like to especially thank Hsiu-Yu Yang and Anna Penzkofer for their support and the numerous insightful discussions.
We used PyTorch (Paszke et al., 2019) to manipulate tensors, Transformers (Wolf et al., 2020) to evaluate the models and NNsight (Fiotto-Kaufman, ) to perform activation interventions. We performed our data analysis using NumPy (Harris et al., 2020), Pandas (pandas development team, 2020; Wes McKinney, 2010), and SciPy (Virtanen et al., 2020). We made our plots using Matplotlib (Hunter, 2007) and seaborn (Waskom, 2021).
Appendix A Appendix
A.1 Experimental Setup
A.1.1 BigToM
BigToM (Gandhi et al., 2023) is constructed using GPT-4 (Achiam et al., 2023) to populate causal templates and combine elements from these templates. Each causal template is set up with a context and a description of the protagonist (e.g. “Noor is working as a barista […]”), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”). The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”). The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs – true or false. Gandhi et al. (2023) generated 200 templates and extracted 25 conditions from each template, resulting in 5,000 test samples. In this work, following Zhu et al. (2024) and Gandhi et al. (2023), we focused on the 6 most important conditions, corresponding to true and false beliefs on the following three tasks:
  • Forward Belief: given the protagonist’s percepts of the causal event, infer their belief: $P(\mathrm{belief}|\mathrm{percept})$.
  • Forward Action: infer the protagonist’s action given their desire and percepts of the causal event. Before inferring the action, one would need to first implicitly infer the protagonist’s belief: $\sum_{\mathrm{belief}}P(\mathrm{action}|\mathrm{percept},\mathrm{belief},\mathrm{desire})$.
  • Backward Belief: infer the protagonist’s belief from observed actions. This requires to first implicitly infer the protagonist’s percepts: $\sum_{\mathrm{percepts}}P(\mathrm{belief}|\mathrm{action},\mathrm{percept},\mathrm{desire})$.
The dataset was released under the MIT license and can be accessed at https://github.com/cicl-stanford/procedural-evals-tom. We report one example for each task in the boxes below, where the text defining true belief or false belief task is shown in blue and red, respectively.
Example of Forward Belief Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk. Noor does not see her coworker swapping the milk.
Does Noor believe the milk pitcher contains oat milk or almond milk?
a) Noor believes the milk pitcher contains oat milk.
b) Noor believes the milk pitcher contains almond milk.
Example of Forward Action Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk. Noor does not see her coworker swapping the milk.
What will Noor do?
a) Noor will make the cappuccino using the milk in the pitcher.
b) Noor will open the fridge once again to take out the oat milk and replace the almond milk with oat milk.
Example of Backward Belief Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor opens the fridge again and reaches for the oat milk. Noor makes the cappuccino using the milk in the pitcher.
Does Noor believe the milk pitcher contains oat milk or almond milk?
a) Noor believes the milk pitcher contains oat milk.
b) Noor believes the milk pitcher contains almond milk.
A.1.2 Linear Probes
Our probing approach is illustrated in Figure 5. We cache activations both at the attention and residual stream level. In our experiments, we use residual stream activations. To perform ITI and compare it to CAA, we also cache attention heads’ activations. We trained the probes using the L-BFGS solver (Liu & Nocedal, 1989) with L2 penalty with inverse of regularisation strength $10$ for a maximum of $1000$ iterations. We use zero as random seed.
Figure 5: Given a tokenised input, we cache the internal activations for all attention heads $h_{i}$, $i=0,\dots,H-1$, and residual streams. In our experiments, we use residual stream activations.
A.1.3 Language models
A detailed summary of the models we use in this work is shown in Table 2. Pythia was released under the Apache 2.0 license. Llama-2 is licensed by Meta for both researchers and commercial entities (Touvron et al., 2023). For all the models, we set the temperature to zero.
Table 2: The 12 models used in this work. The checkmark indicates we additionally study the fine-tuned (Chat) version of the model.
LM Size + Chat Tokens $d_{model}$ Layers
Llama-2 (Touvron et al., 2023) 7B 2T 4096 32
13B 2T 5120 40
70B 2T 8192 80
Pythia (Biderman et al., 2023) 70M 300B 512 6
410M 300B 1024 24
1B 300B 2048 16
6.9B 300B 4096 32
12B 300B 5120 40
A.1.4 Examples of Prompt Variations
Default prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: Noor believes the milk pitcher contains oat milk.
Random prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: 장士 decided [_ countries sections behoSOUR gminy bef Noor believes the milk pitcher contains oat milk.
Misleading prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: Noor believes the milk pitcher contains oat milk.
Belief: The water valve is closed.
Time specification prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: In the end, Noor believes the milk pitcher contains oat milk.
Initial belief prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. Noor believes that the milk pitcher contains oatmilk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: Noor believes the milk pitcher contains oat milk.
A.2 Model Size and Fine-tuning
To characterise the relationship between probe accuracy and model size we consider the best probe accuracy for every LM, i.e. the highest accuracy among probes $\{g_{l}\}$ trained on $\{a_{l}\}$ for a LM $f$. For Llama-2 base, the best probe accuracy scales logarithmically with model size ($R^{2}=0.98$, Figure 6(b)), whereas for fine-tuned models it scales linearly ($R=1.0$, cf. Figure 6(c)). For Pythia base, the best probe accuracy also scales logarithmically with model size ($R^{2}=0.96$, Figure 6(d)).
Figure 6: To characterise the relationship between probe accuracy and model size we consider the best probe accuracy for every LM, i.e. the highest accuracy among probes $\{g_{l}\}$ trained on $\{a_{l}\}$ for a LM $f$. (a) Best accuracy for Llama-2 models of different size. Numbers on the vertical dotted lines indicate the gain in accuracy between base and fine-tuned model of the same size. (b) Logarithmic fit for Llama-2 base. (c) Linear fit for Llama-2 fine-tuned (chat). (d) Logarithmic fit for Pythia base.
A.3 Sensitivity to Prompting
Accuracy on oracle belief probing for different prompt variations are reported in Figure 7.
Figure 7: Sensitivity of protagonist belief probing accuracy to different prompt variations.
A.4 Dimensionality Reduction
Oracle probe accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components are shown in Figure 8.
Figure 8: (Oracle) To investigate potential memorisation in the probes, we compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components. For Llama2: All(7b) = $4096$, All(13b) = $5120$, All(70b) = $8192$. For Pythia: All(70m) = $512$, All(410m) = $1024$, All(1b) = $2048$, All(6.9b) = $4096$, All(12b) = $5120$.
A.5 Inference-time intervention
Inference-time intervention (Li et al., 2023c, ITI) employs a two-step process. First, it trains a probe for each attention head across all layers of a LM. These probes are evaluated on a validation set, and the top-$k$ heads with the highest accuracy are selected. Subsequently, during inference, ITI steers the activations of these top heads along the directions defined by their corresponding probes. Formally, ITI can be defined as an additional term to the multi-head attention:
$x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\left(\mathrm{Att}_{l}^{h}(P_{l}^{h}x_{l}){\color[rgb]{0.5,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,1}\pgfsys@color@cmyk@stroke{0.50}{1}{0}{0}\pgfsys@color@cmyk@fill{0.50}{1}{0}{0}\ +\ \alpha\sigma_{l}^{h}\theta_{l}^{h}}\right)$
where $x_{l}$ is the residual stream at layer $l$, $H$ is the number of attention heads, $\alpha\in\mathbb{R}^{+}$ is a coefficient, $\sigma_{l}^{h}$ is the standard deviation of activations along the direction identified by the probe trained on attention head $h$ at layer $l$, and $\theta_{l}^{h}$ is zero ofr not-selected attention heads.
A.6 Activation editing hyperparameters
Table 3 reports results obtained on the three BigToM tasks with the hyperparameters used for ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023). We report an example of prompt used for evaluation in the box below.
Prompt used for downstream evaluation (Section 4.4) Answer the questions based on the context. Keep your answer concise, few words are enough, maximum one sentence. Answer as ’Answer:<option>)<answer>’.
Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
Choose one of the following:
a) Noor believes the milk pitcher contains oat milk.
b) Noor believes the milk pitcher contains almond milk.
Answer:
Table 3: Comparison between ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023) on the three BigToM tasks (Gandhi et al., 2023). For ITI, the subscript indicates the value of the coefficient used. For CAA, the subscript indicates first the value of the coefficient used and second the layer at which intervention takes place.
Model Method Forward Belief Forward Action Backward Belief
TB FB TB $\land$ FB TB FB TB $\land$ FB TB FB TB $\land$ FB
Llama-2-7b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $44_{0.0}$ $44_{0.0}$ $44_{0.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$
CAA $66_{2.0,11}$ $71_{1.0,31}$ $54_{2.0,0}$ $66_{2.0,11}$ $57_{2.0,12}$ $54_{2.0,2}$ $60_{2.0,11}$ $74_{1.0,31}$ $54_{2.0,2}$
Llama-2-7b-chat No int. $56$ $56$ $55$ $69$ $55$ $37$ $56$ $56$ $55$
ITI $58_{15.0}$ $58_{15.0}$ $57_{15.0}$ $69_{0.0}$ $55_{0.0}$ $37_{0.0}$ $58_{10.0}$ $60_{10.0}$ $57_{10.0}$
CAA $70_{1.0,11}$ $72_{1.5,10}$ $57_{1.0,1}$ $69_{0.0,0}$ $67_{1.5,11}$ $53_{1.5,12}$ $66_{1.0,11}$ $84_{1.5,10}$ $57_{1.0,0}$
Llama-2-13b No int. $52$ $44$ $35$ $59$ $50$ $37$ $46$ $49$ $33$
ITI $52_{0.0}$ $45_{15.0}$ $35_{0.0}$ $64_{15.0}$ $61_{20.0}$ $46_{20.0}$ $48_{20.0}$ $59_{20.0}$ $42_{20.0}$
CAA $85_{2.0,12}$ $88_{2.0,14}$ $66_{2.0,12}$ $71_{1.5,10}$ $69_{2.0,13}$ $55_{1.0,39}$ $75_{2.0,10}$ $92_{2.0,13}$ $59_{1.5,12}$
Llama-2-13b-chat No int. $84$ $56$ $47$ $78$ $51$ $38$ $72$ $48$ $31$
ITI $84_{0.0}$ $65_{15.0}$ $59_{15.0}$ $78_{0.0}$ $58_{15.0}$ $47_{15.0}$ $72_{0.0}$ $60_{15.0}$ $48_{15.0}$
CAA $97_{1.0,12}$ $94_{1.0,12}$ $91_{1.0,12}$ $80_{1.5,11}$ $71_{1.0,13}$ $54_{1.5,13}$ $97_{1.5,10}$ $94_{1.5,12}$ $87_{1.5,12}$
Llama-2-70b No int. $90$ $87$ $78$ $93$ $52$ $48$ $73$ $53$ $32$
ITI $90_{0.0}$ $90_{20.0}$ $78_{0.0}$ $94_{15.0}$ $55_{20.0}$ $50_{15.0}$ $77_{10.0}$ $58_{15.0}$ $37_{10.0}$
CAA $99_{2.0,16}$ $97_{1.5,19}$ $95_{1.5,18}$ $94_{1.5,2}$ $80_{2.0,19}$ $73_{1.5,18}$ $94_{2.0,18}$ $92_{2.0,19}$ $83_{1.5,19}$
Llama-2-70b-chat No int. $69$ $75$ $56$ $86$ $56$ $52$ $63$ $59$ $52$
ITI $69_{0.0}$ $76_{10.0}$ $59_{10.0}$ $86_{0.0}$ $56_{0.0}$ $52_{0.0}$ $63_{0.0}$ $60_{10.0}$ $54_{10.0}$
CAA $92_{1.5,18}$ $97_{1.5,25}$ $89_{1.5,18}$ $87_{1.5,17}$ $75_{1.0,19}$ $60_{1.0,19}$ $88_{1.5,18}$ $92_{1.0,19}$ $80_{1.5,18}$
Pythia-70m No int. $41$ $41$ $37$ $46$ $45$ $41$ $44$ $41$ $37$
ITI $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$
CAA $62_{1.0,2}$ $56_{1.0,1}$ $54_{1.5,1}$ $59_{1.0,2}$ $60_{1.0,3}$ $58_{1.0,2}$ $63_{1.0,2}$ $56_{1.0,2}$ $54_{1.5,1}$
Pythia-410m No int. $48$ $45$ $45$ $44$ $44$ $44$ $44$ $47$ $44$
ITI $55_{20.0}$ $62_{20.0}$ $52_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $60_{20.0}$ $63_{20.0}$ $56_{20.0}$
CAA $67_{2.0,4}$ $64_{2.0,4}$ $61_{2.0,0}$ $56_{2.0,6}$ $63_{1.5,12}$ $56_{2.0,6}$ $69_{2.0,4}$ $63_{2.0,0}$ $60_{2.0,0}$
Pythia-1b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$
CAA $59_{2.0,8}$ $62_{2.0,5}$ $54_{2.0,0}$ $57_{2.0,4}$ $59_{2.0,10}$ $56_{2.0,4}$ $57_{2.0,3}$ $60_{2.0,5}$ $54_{2.0,0}$
Pythia-6.9b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $45_{20.0}$ $54_{20.0}$ $44_{0.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$
CAA $56_{1.5,12}$ $71_{1.5,9}$ $55_{2.0,23}$ $55_{2.0,4}$ $63_{1.5,11}$ $55_{2.0,4}$ $55_{2.0,23}$ $71_{1.5,9}$ $55_{2.0,23}$
Pythia-6.9b-chat No int. $55$ $54$ $28$ $36$ $64$ $20$ $44$ $67$ $30$
ITI $57_{15.0}$ $54_{0.0}$ $28_{0.0}$ $44_{15.0}$ $71_{15.0}$ $32_{15.0}$ $44_{0.0}$ $67_{0.0}$ $30_{0.0}$
CAA $68_{1.5,15}$ $65_{1.5,12}$ $57_{1.5,11}$ $54_{1.5,10}$ $75_{1.5,5}$ $48_{1.5,10}$ $58_{1.5,15}$ $67_{0.0,0}$ $54_{1.5,10}$
Pythia-12b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
ITI $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$ $54_{20.0}$
CAA $54_{2.0,0}$ $64_{2.0,9}$ $54_{2.0,0}$ $60_{2.0,11}$ $58_{2.0,11}$ $55_{2.0,12}$ $54_{2.0,0}$ $67_{2.0,10}$ $54_{2.0,0}$
A.7 Compute Resources
We ran our experiments on a server running Ubuntu 22.04, equipped with eight NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs.
A.8 Code
Our code will be made public under the MIT licence at https://git.hcics.simtech.uni-stuttgart.de/public-projects/mental-states-in-LMs.

### v3

@@
-Benchmarking Mental State Representations in Language Models
+Brittle Minds, Fixable Activations: Understanding Belief Representations in Language Models
 Abstract
-While numerous works have assessed the generative performance of language models (LMs) on tasks requiring Theory of Mind reasoning, research into the models’ internal representation of mental states remains limited.
-Recent work has used probing to demonstrate that LMs can represent beliefs of themselves and others.
-However, these claims are accompanied by limited evaluation, making it difficult to assess how mental state representations are affected by model design and training choices.
-We report an extensive benchmark with various LM types with different model sizes, fine-tuning approaches, and prompt designs to study the robustness of mental state representations and memorisation issues within the probes.
-Our results show that the quality of models’ internal representations of the beliefs of others increases with model size and, more crucially, with fine-tuning.
-We are the first to study how prompt variations impact probing performance on theory of mind tasks.
-We demonstrate that models’ representations are sensitive to prompt variations, even when such variations should be beneficial.
-Finally, we complement previous activation editing experiments on Theory of Mind tasks and show that it is possible to improve models’ reasoning performance by steering their activations without the need to train any probe.
-Machine Learning, ICML
+Despite growing interest in Theory of Mind (ToM) tasks for evaluating language models (LMs), little is known about how LMs internally represent mental states of self and others.
+Understanding these internal mechanisms is critical – not only to move beyond surface-level performance, but also for model alignment and safety, where subtle misattributions of mental states may go undetected in generated outputs.
+In this work, we present the first systematic investigation of belief representations in LMs by probing models across different scales, training regimens, and prompts – using control tasks to rule out confounds.
+Our experiments provide evidence that both model size and fine-tuning substantially improve LMs’ internal representations of others’ beliefs, which are structured – not mere by-products of spurious correlations – yet brittle to prompt variations.
+Crucially, we show that these representations can be strengthened: targeted edits to model activations can correct wrong ToM inferences.
+Brittle Minds, Fixable Activations:
+Understanding Belief Representations in Language Models
+Matteo Bortoletto  Constantin Ruhdorfer  Lei Shi  Andreas Bulling University of Stuttgart, Germany matteo.bortoletto@vis.uni-stuttgart.de
 1 Introduction
-Modern language models (LMs) trained on next token prediction have demonstrated impressive capabilities, spanning coding, mathematical reasoning, fact verification, and embodied interaction (Wei et al., 2022; Bubeck et al., 2023).
-As these models are designed with the ultimate goal of collaborating with humans, it becomes imperative that they complement these skills with an understanding of humans, in particular their beliefs, emotions, desires, and intentions (Li et al., 2023a).
-Core to this understanding is Theory of Mind (ToM) – the ability to attribute mental states to oneself and others (Premack & Woodruff, 1978).
+Language models (LMs) trained on next token prediction have demonstrated impressive capabilities across various tasks, spanning coding, math, and embodied interaction  (Wei et al., 2022; Bubeck et al., 2023).
+As these models are designed with the ultimate goal of collaborating with humans, it becomes imperative that they complement these skills with an understanding of humans.
+Core to this understanding is Theory of Mind (ToM) – the ability to attribute mental states to oneself and others (Premack and Woodruff, 1978).
 ToM is essential for effective communication and cooperation with other agents, facilitating interaction and learning from feedback and demonstrations (Saha et al., 2023).
-Given its significance, ToM has emerged as a critical milestone in artificial intelligence (AI) and an important capability when evaluating cutting-edge LMs (Bubeck et al., 2023).
-Interest in LMs’ generative performance on tasks requiring ToM reasoning has resulted in a wide variety of benchmark datasets, typically involving question-answering tasks (Le et al., 2019; Gandhi et al., 2023; Kim et al., 2023; He et al., 2023; Tan et al., 2024; Xu et al., 2024).
-Despite showing improved performance on ToM benchmarks compared to earlier models, modern LMs are still far from perfect (Sap et al., 2022).
-Text generated by LMs often contains errors that limit their performance on ToM tasks (Martindale et al., 2019).
-Previous work has shown that it is sometimes possible to still obtain correct predictions by probing LMs’ internal representations (Li et al., 2021; Liu et al., 2023b; Gurnee et al., 2023).
-In particular, Zhu et al.
-(2024) have shown that LMs, when prompted with a story and a belief statement, can represent beliefs from their own perspective and, to a lesser extent, from the perspective of a character in the story.
-Their work is an important first step towards understanding how LMs represent mental states, but it is limited in the number of models and settings studied, leaving many questions unanswered.
-Building and extending on (Zhu et al., 2024), we benchmark mental state representations of self and others in language models through extensive experiments of different LM families, model sizes, and fine-tuning approaches.
-Specifically, we design a set of experiments to address the following research questions: RQ1.
-What is the relation between model size and probing accuracy?
-RQ2.
-Does fine-tuning with instruction-tuning (Wei et al., 2021) and/or reinforcement learning from human feedback (Christiano et al., 2017; Ouyang et al., 2022, RLHF) have an effect on probing accuracy?
-RQ3.
-Are models’ internal representations of beliefs sensitive to prompt variations?
-RQ4.
-Is there a risk of probes memorising training data due to the large dimensionality of LM representations?
-RQ5.
-Can we enhance LMs’ performance by editing their activations without training dedicated probes?
-To answer RQ1, we perform probing experiments on two families of LMs, Llama-2 (Touvron et al., 2023), and Pythia (Biderman et al., 2023), ranging from models with 70 million to 70 billion parameters.
-To address RQ2, we compare the probing performance of models pre-trained solely on next token prediction with models that have been fine-tuned using instruction-tuning and/or RLHF.
-Our experiments reveal that probing accuracy on the beliefs of others increases with model size and, more crucially, with fine-tuning.
-To answer RQ3, we repeat our probing experiments using different variations of the prompt used by Zhu et al..
-Specifically, we consider two variations that are expected to negatively impact LMs’ representations (Random, Misleading), and two that are supposed to have a positive influence (Time Specification, Initial Belief).
-By conducting these experiments, our work is the first to explore the sensitivity of LMs’ representations to prompting in the context of ToM.
-Our findings demonstrate that models’ representations are sensitive to prompt variations, even when such variations should be beneficial.
-To address RQ4, we compare our trained probes with a second set of probes trained only on the representations’ first top $k$ principal components.
-This requires learning much fewer parameters and eliminates any possible memorisation issue.
-We find no strong evidence of memorisation in the probes, as it is possible to recover most of the accuracy by training probes on a small subset of principal components of models’ representations.
-We formulate RQ5 as a follow-up question to Zhu et al.
-(2024) who found that probes trained to predict beliefs can be used to steer models’ activation using inference-time intervention (Li et al., 2023c, ITI) to improve LMs’ downstream performance on ToM tasks.
-In contrast, we show that by using contrastive activation addition (Rimsky et al., 2023, CAA), we can steer models’ activations without the need to train any probe and, in a more generalisable way, obtain significant performance improvements across different ToM tasks.
+Given its significance, computational ToM has emerged as a key capability when evaluating cutting-edge LMs (Ma et al., 2023; Shapira et al., 2024; Chen et al., 2025).
+Figure 1: ToM tasks are challenging for LMs, but correct predictions can sometimes be recovered by probing their internal representations.
+We study how internal representations of beliefs of self and others emerge in 12 LMs, and show that these representations are structured yet brittle to prompts, and can be strengthened with a steering vector to fix incorrect ToM inferences.
+Despite the improved performance on ToM benchmarks compared to earlier models, modern LMs are still far from perfect (Sap et al., 2022).
+Text generated by LMs often contains errors that limit their performance on ToM tasks.
+Zhu et al.
+(2024) showed that probing LMs’ internal representations can sometimes recover correct belief inferences, with models like Mistral-7B-Instruct Jiang et al.
+(2023) and DeepSeek-7B-Chat Bi et al.
+(2024) capturing beliefs from both their own and others’ perspectives.
+While promising, this remains a preliminary step: it examines only single-sized, fine-tuned models, leaves possible confounds uncontrolled, and ignores how subtle changes in prompting affect belief representations.
+As a result, we still lack a clear understanding of how internal belief representations differ across models, whether they reflect true ToM or spurious patterns, and how robust they are to prompts.
+To address these gaps, we pose four key research questions and present evidence for each. We begin by studying how emergence scales across models:
+RQ1
+Do internal belief representations emerge similarly in different LMs, and are they affected by model size and training regime?
+Finding training regimes or scales that are more conducive to belief reasoning can guide future model development toward more reliable ToM behaviour.
+However, it is also crucial to verify if representations are structured, indicating genuine modelling of mental states, or spurious:
+RQ2
+Are LMs’ internal belief representations structured or the result of spurious correlations?
+This distinction is essential for determining if representations reflect a genuine understanding of beliefs or only exploit statistical patterns that happen to correlate with correct answers in the training data.
+This is also crucial for alignment and safety, as misaligned mental state attributions may not appear overtly in text – leading to false signals of understanding.
+Equally important is that models can maintain robust belief attributions:
+RQ3
+Are LMs’ internal belief representations robust?
+Fragile representations may break under slight variations, leading to inconsistent or unsafe behaviour in real-world applications involving social reasoning or user interaction.
+Strengthening these representations, then, offers a promising path toward improving their reliability:
+RQ4
+Can we strengthen LMs’ internal belief representations to improve their performance?
+To answer these research questions, we perform probing and activation editing experiments using 12 LMs (Figure 1).
+We first compare base models with those fine-tuned via SFT and/or RLHF Ouyang et al.
+(2022)(RQ1), finding that belief representations emerge in consistent patterns across models, improve with model size, and – especially in smaller models – benefit significantly from fine-tuning.
+To provide evidence that LMs’ belief representations are structured (RQ2), we show that (1) probes trained on randomly permuted labels perform at chance – confirming selectivity, and (2) probes trained on top-$k$ principal components still recover most accuracy for $k\ll d_{model}$.
+Next, we test robustness (RQ3) using varied prompts.
+Surprisingly, semantically neutral changes can reduce accuracy, revealing that representations of others’ beliefs are brittle to prompts.
+However, we show that it is possible to strengthen models’ representation by using contrastive activation addition (Rimsky et al., 2023, CAA), obtaining significant performance improvements across different ToM tasks (RQ4).
 In summary, our work makes the following contributions:
   1. 1.
-    We report extensive probing experiments with various types of LMs with different model sizes and fine-tuning approaches, showing that the quality of models’ internal representations of the beliefs of others increases with model size and, more crucially, fine-tuning.
+    We provide extensive probing experiments across 12 LMs, suggesting that representations of others’ beliefs improve with size and fine-tuning, and that these representations are structured yet brittle to prompt variations.
   2. 2.
-    We are the first to study how prompt variations impact belief probing performance, showing that models’ representations are sensitive to prompt variations, even when such variations should be beneficial.
-  3. 3.
-    We show that by using contrastive activation addition it is possible to improve models’ reasoning performance by steering their activations without the need to train any probe.
+    We show that we can strengthen models’ representations by using contrastive activation addition and improve their ToM performance.
 2 Related Work
 Machine Theory of Mind
-Theory of Mind (ToM) has been studied in cognitive science and psychology for decades (Gurney et al., 2021).
-Mirroring efforts to understand ToM in humans, an increasing number of works in the computational sciences have investigated means to equip AI with similar capabilities.
-Previously proposed models that aim to implement a machine ToM have been based on partially observable Markov decision processes (POMDP) (Doshi et al., 2010; Han & Gmytrasiewicz, 2018), Bayesian methods (Baker et al., 2011, 2017) and deep learning methods (Rabinowitz et al., 2018; Bara et al., 2021; Wang et al., 2022; Duan et al., 2022; Liu et al., 2023a; Bortoletto et al., 2024b, a).
-Recent advances in LMs have sparked interest in evaluating their ToM capabilities.
-Various benchmarks have been proposed, aiming to measure LMs’ ability to understand and reason about the beliefs, goals, and intentions of others (Le et al., 2019; He et al., 2023; Kim et al., 2023; Gandhi et al., 2023; Xu et al., 2024; Tan et al., 2024).
-Additionally, efforts have been made to enhance LMs’ ToM through prompting techniques (Zhou et al., 2023b; Moghaddam & Honey, 2023; Wilf et al., 2023).
-A new direction of research explores LMs’ internal representation of mental states.
+Theory of mind has been studied in AI for more than a decade (Baker et al., 2009; Rabinowitz et al., 2018; Bara et al., 2021; Bortoletto et al., 2024a, b, c).
+Various benchmarks have been proposed, aiming to measure LMs’ ability to understand and reason about the beliefs, goals, and intentions of others (Le et al., 2019; He et al., 2023; Kim et al., 2023; Gandhi et al., 2023; Xu et al., 2024; Tan et al., 2024; Sclar et al., 2023; Ma et al., 2023; Wu et al., 2023).
+Additionally, efforts have been made to enhance LMs’ ToM through prompting techniques (Zhou et al., 2023b; Moghaddam and Honey, 2023; Wilf et al., 2023).
+Our work dives deeper into LMs’ internal belief representations, offering a broader insight into these mechanisms that go beyond surface-level performance.
+Probing Neural Representations
+Initially proposed by Alain and Bengio (2017), probing is a widely used method for determining if models represent particular features or concepts.
+In the realm of LMs, numerous works used probing to demonstrate that these models acquire rich linguistic representations – spanning semantic concepts such as syntactic categories, dependency relations, co-reference, and word meaning (Conneau et al., 2018; Tenney et al., 2018, 2019; Rogers et al., 2021; Li et al., 2021; Hernandez and Andreas, 2021; Marks and Tegmark, 2023; Liu et al., 2023).
+A separate line of work explored if LMs possess a world model (Li et al., 2021; Abdou et al., 2021; Patel and Pavlick, 2022; Li et al., 2023a; Nanda et al., 2023).
+An emergent line of work that is relevant to our work used probing to explore if LMs have agent models, for example, if they can represent beliefs of self and others (Zhu et al., 2024; Bortoletto et al., 2024a).
+In this work, we contribute with extensive experiments that characterise models’ representations of beliefs along different axes: emergence, structure, robustness, and steerability.
+Prompt Analysis
+Previous work has shown that LMs are vulnerable to prompt alterations like token deletion or reordering (Ishibashi et al., 2023), biased or toxic prompts (Shaikh et al., 2023) and similarity to training data (Razeghi et al., 2022).
+Other works have shown the importance of input-output format (Min et al., 2022) and of demonstration example ordering for few-shot performance (Zhao et al., 2021; Lu et al., 2022; Zhou et al., 2023a).
+In this work, we shift our focus from analysing how sensitive model outputs are to how model representations change Gurnee and Tegmark (2024).
+In particular, we explore for the first time the effect of prompt variations on how models internally represent mental states.
+Activation Editing
+Activation editing has emerged as a way to influence model behaviour without any additional fine-tuning (Li et al., 2023a; Hernandez et al., 2023).
+One notable method in this domain is inference-time intervention (Li et al., 2023b, ITI), which involves training linear probes on contrastive question-answering datasets to identify “truthful” attention heads and then shifting their activations during inference along the identified truthful directions.
+In contrast, activation addition (Turner et al., 2023, AA) and contrastive activation addition (Rimsky et al., 2023, CAA) generate steering vectors by only using LMs’ activations.
 Zhu et al.
-(2024) demonstrated that LMs linearly encode beliefs from different agents’ perspectives, and manipulating these representations can enhance ToM task performance.
-While Zhu et al.’s work is a crucial initial step, our work dives deeper into LMs’ internal belief representations, offering a broader insight into these mechanisms.
-Probing Neural Representations
-Initially proposed by Alain & Bengio (2017), probing has emerged as a common method for determining if models represent particular features or concepts.
-In the realm of LMs, numerous works used probing to demonstrate that these models acquire rich linguistic representations.
-These representations span syntactic and semantic concepts such as syntactic categories, dependency relations, co-reference, and word meaning (Conneau et al., 2018; Tenney et al., 2018, 2019; Rogers et al., 2021; Li et al., 2021; Hernandez & Andreas, 2021; Marks & Tegmark, 2023; Liu et al., 2023b).
-A separate line of work explored if and how LMs represent the world, i.e., whether they possess a world model.
-Li et al.
-(2021) showed that LMs track the states of entities within a context.
-Other works showed that LMs exhibit representations reflecting non-linguistic concepts in the world, which LMs have never observed (Abdou et al., 2021; Patel & Pavlick, 2022; Li et al., 2023b; Nanda et al., 2023).
-An emergent line of work that is particularly relevant to our work used probing to explore if LMs have agent models, for example, if they can represent beliefs of self and others (Zhu et al., 2024; Bortoletto et al., 2024a).
-While representing an important first step towards understanding the internals of Theory of Mind in LMs, experiments in (Zhu et al., 2024) are limited in settings and models considered.
-In this work, we contribute with extensive experiments that employ a wider variety of LMs and a wider range of settings.
-Furthermore, we also explore possible memorisation issues in the probes.
-Prompt Analysis
-Research on prompt robustness in LMs is still in its infancy but has quickly sparked much interest.
-On one hand, previous work has shown that LMs are vulnerable to prompt alterations like token deletion or reordering (Ishibashi et al., 2023), biased or toxic prompts (Shaikh et al., 2023) and similarity to training data (Razeghi et al., 2022).
-On the other hand, instruction-tuned models have proved to be more robust against prompt variation, even when using misleading instructions (Webson & Pavlick, 2022).
-Other works have shown the importance of input-output format (Min et al., 2022) and of demonstration example ordering for few-shot performance (Zhao et al., 2021; Lu et al., 2022; Zhou et al., 2023a).
-In this work, we shift our focus from analysing how sensitive model outputs are to how model representations change.
-Our work, along with (Gurnee et al., 2023), is one of the first to explore how prompt design affects how accurately models represent concepts.
-In particular, Gurnee et al.
-(2023) have studied whether LMs’ representations of space and time are robust to prompt variations.
-In stark contrast, we explore for the first time the effect of prompt variations on how models represent mental states internally.
-Activation Editing
-Recent advancements in NLP have introduced innovative techniques for controlling and manipulating text generation models.
-While weight editing proposed to modify models’ weights (Meng et al., 2022; Ilharco et al., 2022; Orgad et al., 2023), activation editing has emerged as an alternative way to influence model behaviour without any additional fine-tuning (Li et al., 2023b; Hernandez et al., 2023).
-This approach involves manipulating the internal representations of models to direct their outputs towards desired outcomes.
-One notable method in this domain is inference-time intervention (Li et al., 2023c, ITI), which has been proposed to enhance truthfulness in LMs.
-ITI involves training linear probes on contrastive question-answering datasets to identify “truthful” attention heads and then shifting attention head activations during inference along the identified truthful directions.
-In contrast, activation addition (Turner et al., 2023, AA) and contrastive activation addition (Rimsky et al., 2023, CAA) offer ways to generate steering vectors by only using LMs’ intermediate activations.
-Zhu et al.
-have used ITI to show that it is possible to manipulate LMs’ internal representations of mental states (Zhu et al., 2024).
-In this work, we show that using CAA can further improve LMs’ ToM capabilities without the necessity of training any probe.
-Remarkably, CAA operates at the residual stream level, eliminating the need for a fine-grained search over attention heads.
+has used ITI to show that it is possible to manipulate LMs’ internal representations of mental states.
+In this work, we show that using CAA can further improve LMs’ ToM capabilities while eliminating the need for a fine-grained search over attention heads.
 3 Experimental Setup
 3.1 Probing
+We linearly decode belief status from the perspective of different agents by using probing (Alain and Bengio, 2017).
+Probing involves localising specific concepts in a neural model by training a simple classifier (called a probe) on model activations to predict a target label associated with the input data.
+To provide a formal definition, we adopt a similar notation to the one introduced in (Belinkov, 2022).
+Consider an original model $f:x\mapsto\hat{y}$ that is trained on a dataset $\mathcal{D}^{O}=\{x^{(i)},y^{(i)}\}$ to map input $x$ to output $\hat{y}$.
+Model performance is evaluated by some measure, denoted $\textsc{Perf}(f,\mathcal{D}^{O})$.
+A probe $g_{l}:f_{l}(x)\mapsto\hat{z}$ maps intermediate representations of $x$ in $f$ at layer $l$ to some property $\hat{z}$, which is the label of interest.
+The probe $g_{l}$ is trained on a probing dataset $\mathcal{D}^{P}=\{x^{(i)},z^{(i)}\}$ and evaluated using some performance measure $\textsc{Perf}(g_{l},f,\mathcal{D}^{O},\mathcal{D}^{P})$.
+In our case, $f$ is an autoregressive language model that, given a sequence of tokens $x$, outputs a probability distribution over the token vocabulary to predict the next token in the sequence.
+Our probe is a logistic regression model $g_{l}:\hat{z}=Wa_{l}+b$ trained on neural activations $f_{l}(x)=a_{l}$ to predict binary belief labels $y=\{0,1\}$.
+3.2 Dataset
+We use BigToM (Gandhi et al., 2023), a question-answering dataset constructed by populating causal templates and combining elements from these templates.
+Each causal template is set up with a context and a description of the protagonist (e.g.
+“Noor is working as a barista […]”, see Story in Figure 2), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”).
+The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”).
+The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs.
+Similar to (Zhu et al., 2024), we focus on the Forward Belief setting in which models have to infer the belief of the protagonist given the percepts of the causal event, $P(\mathrm{belief}|\mathrm{percepts})$.
+We report additional details in Section A.1.1
 Figure 2: Example of false belief from our probing datasets.
 The labels $z_{p}$ and $z_{o}$ correspond to $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$, respectively.
-By manipulating the protagonist’s percepts after the causal event we obtain two scenarios: true belief and false belief.
-In line with previous work (Zhu et al., 2024) we linearly decode belief status from the perspective of different agents by using probing (Alain & Bengio, 2017).
-Probing involves localising specific concepts in a neural model by training a simple classifier (called a probe) on model activations to predict a target label associated with the input data.
-To provide a formal definition, we adopt a similar notation to the one introduced in (Belinkov, 2022).
-Let us define an original model $f:x\mapsto\hat{y}$ that is trained on a dataset $\mathcal{D}^{O}=\{x^{(i)},y^{(i)}\}$ to map input $x$ to output $\hat{y}$.
-Model performance is evaluated by some measure, denoted $\textsc{Perf}(f,\mathcal{D}^{O})$.
-A probe $g:f_{l}(x)\mapsto\hat{z}$ maps intermediate representations of $x$ in $f$ at layer $l$ to some property $\hat{z}$, which is the label of interest.
-The probe $g$ is trained on a probing dataset $\mathcal{D}^{P}=\{x^{(i)},z^{(i)}\}$ and evaluated using some performance measure $\textsc{Perf}(g,f,\mathcal{D}^{O},\mathcal{D}^{P})$.
-In our case, $f$ is an autoregressive language model that given a sequence of tokens $x$ outputs a probability distribution over the token vocabulary to predict the next token in the sequence.
-Our probe is a logistic regression model $\hat{z}=\sigma(Wa_{l}+b)$ trained on neural activations $f_{l}(x)=a_{l}$ to predict belief labels $y=\{0,1\}$.
-3.2 Dataset
-Following Zhu et al.
-(2024) we use the BigToM benchmark (Gandhi et al., 2023).
-BigToM is constructed using GPT-4 (Achiam et al., 2023) to populate causal templates and combine elements from these templates.
-Each causal template is set up with a context and a description of the protagonist (e.g.
-“Noor is working as a barista […]”), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”).
-The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”).
-The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs.
-In this work, we focus on the Forward Belief setting proposed by (Zhu et al., 2024) in which models have to infer the belief of the protagonist given the percepts of the causal event, $P(\mathrm{belief}|\mathrm{percepts})$.
-We report additional details in Section A.1.1
+By manipulating the protagonist’s percepts after the causal event, we obtain two scenarios: true belief and false belief.
 Probing Datasets
@@
 $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ are built by pairing each story in BigToM with a belief statement, as shown in Figure 2.
-After prompting the model with a story-belief pair we cache the residual stream activations at the final token position for all residual streams (Figure 5).
+After prompting the model with a story-belief pair $x$ we cache the residual stream activations $f_{l}(x)$ at the final token position for all residual streams (see Figure 6).
 3.3 Models
-Zhu et al.
-(2024) have used two models for their experiments: Mistral-7B-Instruct (Jiang et al., 2023) and DeepSeek-7B-Chat (Bi et al., 2024) – both being the same size and fine-tuned.
-In contrast, we study two families of LMs that offer us options in model sizes and fine-tuning: Pythia (Biderman et al., 2023) and Llama-2 (Touvron et al., 2023).
-While Llama-2 offers “chat” versions fine-tuned using supervised learning and RLHF, Pythia’s open-source training set (Gao et al., 2020) ensures that there is no data leakage [^1] Llama-2 was released later than BigToM..
-Additionally, we consider a version of Pythia-6.9B fine-tuned on a mixture of open-source instruction datasets (Wang et al., 2024), which we refer to as Pythia-6.9B-chat.
-[^2] https://huggingface.co/allenai/open-instruct-pythia-6.9b-tulu A summary of the models we study is reported in Table 2.
+We study two families of LMs that offer us options in model sizes and fine-tuning: Pythia (Biderman et al., 2023) and Llama-2 (Touvron et al., 2023) – for a total of 12 models.
+While Llama-2 offers “chat” versions first trained with SFT and then RLHF, Pythia’s open-source training set (Gao et al., 2020) ensures that there is no data leakage [^1] Llama-2 was released later than BigToM..
+Additionally, we consider a SFT version of Pythia-6.9B trained on open-source instruction datasets (Wang et al., 2024), which we refer to as Pythia-6.9B-chat.
+[^2] https://huggingface.co/allenai/open-instruct-pythia-6.9b-tulu We provide model details in Table 2.
 3.4 Probing Experiments
-We aim to contribute to understanding how LMs represent beliefs of self and others by proposing a set of extensive probing experiments across LMs that differ in architecture, size, and fine-tuning approach.
-Our approach is generally similar to the one used by (Zhu et al., 2024), but we make a different operational choice: While (Zhu et al., 2024) train probes on each attention head for every layer, we train probes on the residual stream for every layer.
-We opted to use the residual stream as it integrates information from both the attention and feed-forward components, potentially encoding richer representations.
+To study how LMs represent beliefs of self and others, we propose a set of extensive probing experiments across LMs that differ in architecture, size, and fine-tuning regime.
+We train probes on the residual stream, as it integrates information from both the attention and feed-forward components, potentially encoding richer representations.
 Additionally, since the residual activations directly contribute to the final output predictions, probing them may better align with understanding the model’s behaviour for downstream tasks.
-Model Size and Fine-tuning
-We first report experiments to better understand the effect of model size and fine-tuning on belief probing accuracy.
-Specifically, we ask the following questions: Is there a relation between model size and probing accuracy?
-(RQ1) Does fine-tuning an LM with instruction-tuning or RLHF have an effect on probing accuracy?
-(RQ2) To answer these questions we performed the same probing experiment across all our models and compared the results.
-Figure 2: Belief probing accuracy across models with different architecture, size and fine-tuning.
-Figure 3: Sensitivity of protagonist belief probing accuracy to different prompt variations.
-Figure 4: To investigate potential memorisation in the probes, we compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components.
-For Llama2: All(7b) = $4096$, All(13b) = $5120$, All(70b) = $8192$.
-For Pythia: All(70m) = $512$, All(410m) = $1024$, All(1b) = $2048$, All(6.9b) = $4096$, All(12b) = $5120$.
-We report results for Protagonist beliefs.
-Results for Oracle are shown in Figure 8.
-Table 1: Comparison of the effects of ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023) activation editing methods on three tasks from BigToM (Gandhi et al., 2023).
-TB denotes a true belief task, whereas FB denotes a false belief task.
-The numbers represent accuracy scores, with the difference in performance compared to no intervention (No int.) indicated as subscripts (ITI -- No int.
-and CAA -- No int.).
-An asterisk (*) denotes a statistically significant difference from No int.
-based on a McNemar’s test (McNemar, 1947) with $p<0.05$.
-Model Method Forward Belief Forward Action Backward Belief
-TB FB TB $\land$ FB TB FB TB $\land$ FB TB FB TB $\land$ FB
-Llama-2-7b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
-ITI $44_{+0}$ $44_{+0}$ $44_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-CAA $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $74_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+30}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-Llama-2-7b-chat No int. $56$ $56$ $55$ $69$ $55$ $37$ $56$ $56$ $55$
-ITI $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $69_{+0}$ $55_{+0}$ $37_{+0}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$
-CAA $70_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}$ $72_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $69_{+0}$ $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $53_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $84_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}^{*}$
-Llama-2-13b No int. $52$ $44$ $35$ $59$ $50$ $37$ $46$ $49$ $33$
-ITI $52_{+0}$ $45_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $35_{+0}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$ $61_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $46_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $42_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$
-CAA $85_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+33}}^{*}$ $88_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+44}}^{*}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+31}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $69_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}^{*}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+29}}^{*}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+43}}^{*}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+26}}^{*}$
-Llama-2-13b-chat No int. $84$ $56$ $47$ $78$ $51$ $38$ $72$ $48$ $31$
-ITI $84_{+0}$ $65_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $78_{+0}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $47_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}^{*}$ $72_{+0}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}$
-CAA $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+38}}^{*}$ $91_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+44}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+20}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+46}}^{*}$ $87_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+56}}^{*}$
-Llama-2-70b No int. $90$ $87$ $78$ $93$ $52$ $48$ $73$ $53$ $32$
-ITI $90_{+0}$ $90_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $78_{+0}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $50_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $77_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+4}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$ $37_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$
-CAA $99_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $95_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}^{*}$ $73_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+21}}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+39}}^{*}$ $83_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+51}}^{*}$
-Llama-2-70b-chat No int. $69$ $75$ $56$ $86$ $56$ $52$ $63$ $59$ $52$
-ITI $69_{+0}$ $76_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $86_{+0}$ $56_{+0}$ $52_{+0}$ $63_{+0}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$
-CAA $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+23}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+22}}^{*}$ $89_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+32}}^{*}$ $87_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}^{*}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+8}}^{*}$ $88_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+33}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}$
-Pythia-70m No int. $41$ $41$ $37$ $46$ $45$ $41$ $44$ $41$ $37$
-ITI $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+8}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}$
-CAA $62_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+21}}^{*}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$
-Pythia-410m No int. $48$ $45$ $45$ $44$ $44$ $44$ $44$ $47$ $44$
-ITI $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $62_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $52_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$
-CAA $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $61_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $69_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$
-Pythia-1b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
-ITI $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-CAA $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}^{*}$ $62_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+15}}$ $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-Pythia-6.9b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
-ITI $45_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $44_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-CAA $56_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $63_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+27}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$
-Pythia-6.9b-chat No int. $55$ $54$ $28$ $36$ $64$ $20$ $44$ $67$ $30$
-ITI $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $54_{+0}$ $28_{+0}$ $44_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+8}}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $32_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $44_{+0}$ $67_{+0}$ $30_{+0}^{*}$
-CAA $68_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}$ $65_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $57_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+29}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}^{*}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}^{*}$ $67_{+0}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+24}}^{*}$
-Pythia-12b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
-ITI $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-CAA $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+20}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+14}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $67_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+23}}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$
-Sensitivity to Prompting
-By using a single prompt design, previous work left the impact of prompt design on probing accuracy unclear (Zhu et al., 2024).
-Our second set of experiments aims to explore how belief representations are sensitive to different prompts.
-Research on prompt robustness in language models is still in its infancy and focused mainly on revealing vulnerability to prompt alternations on downstream performance (Min et al., 2022; Ishibashi et al., 2023; Shaikh et al., 2023; Leidinger et al., 2023).
-In contrast, we study how the input influences models’ representations by asking: Are models’ internal belief representations robust to prompt variations?
-(RQ3) To answer this question we define four prompt variations:
-  • Random: Following (Gurnee & Tegmark, 2024), we add 10 random tokens to the belief statement.
+Control Tasks
+Depending on the model dimension, the probes we train have a significant number of learnable parameters – up to $16,385$ for Llama-2-70B.
+This raises the concern that probes might learn to rely on irrelevant patterns in the data instead of capturing meaningful relationships.
+To account for the potential confounding effect of hidden state size, we include two controls.
+First, following Hewitt and Liang (2019), we train and evaluate probes on a version of $\mathcal{D}^{P}_{p}$ with randomly permuted labels – thus removing real input-label relationships.
+If a probe still performs well on the permuted data, this suggests it may be exploiting superficial correlations rather than capturing genuine structure.
+Second, we effectively reduce the number of learnable parameters in the probes by projecting $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ onto their $k$ largest principal components using PCA before training.
+This minimises the risk of the probes relying on spurious patterns in the data.
+Robustness Tests
+Previous work left the impact of prompting on belief probing accuracy unexplored.
+Our second set of experiments aims to study whether belief representations are robust to different prompts.
+Research on prompt robustness in language models focused mainly on revealing vulnerability to prompt alterations on downstream performance (Min et al., 2022; Ishibashi et al., 2023; Shaikh et al., 2023; Leidinger et al., 2023; Sclar et al., 2024).
+In contrast, we study how different prompt alterations influence probing performance, i.e.
+models’ internal representations.
+Unlike model outputs that are shaped by decoding strategies, which act as confounders, models’ activations are more abstract and offer a better lens into how robust or brittle internal representations are.
+We define four prompt variations:
+  • Random: Following Gurnee and Tegmark (2024), we add 10 random tokens to the belief statement.
   • Misleading: Each story is followed by two belief statements, one pertinent to the story and one randomly chosen from another.
   • Time Specification: The prompt specifies that the belief statement refers to the end of the story.
-We study this variation because some belief statements can be true (false) at the story’s beginning but false (true) at the end.
-For example, consider the story in Figure 1: if Noor does not witness the swap, in the end, she will believe the pitcher contains almond milk ($y_{p}=\mathrm{True}$).
-However, if the same belief is referred to at the beginning of the story, then it is false ($y_{p}=\mathrm{False}$).
+We include this variation because some belief statements can be true (false) at the story’s beginning but false (true) at the end.
+For example, consider the story in Figure 2: if Noor does not witness the swap, in the end, she will believe the pitcher contains almond milk ($z_{p}=\mathrm{True}$).
+However, if the same belief is referred to the beginning of the story, then it is false ($z_{p}=\mathrm{False}$).
   • Initial Belief: We explicitly reveal the protagonist’s initial belief (e.g. “Noor believes that the pitcher contains oat milk”) in the story to test whether it biases the representations of LMs.
-While all maintaining conceptual and semantic parity with the Original prompt used in (Zhu et al., 2024), Random and Misleading are expected to negatively impact LMs’ representations, while Time Specification and Initial Belief are supposed to have a positive influence.
-Robust representations of mental states should exhibit minimal sensitivity to these alterations.
+While all maintain conceptual and semantic parity with the Original prompt used in (Zhu et al., 2024), Random and Misleading are expected to negatively impact LMs’ representations, while Time Specification and Initial Belief are supposed to have a positive influence.
+Robust representations of beliefs should exhibit minimal sensitivity to these alterations.
 Our experiments compare probe accuracy across different model sizes, fine-tuning, and prompt variations.
 Examples of prompts are reported in Section A.1.4.
-Memorisation
-Although linear, our probes possess many learnable parameters – up to $16,385$ for Llama-2-70B.
-In principle, this allows them to engage in significant memorisation (Alain & Bengio, 2017).
-Our final set of probing experiments answers the following question: Are the probes memorising their training data?
-(RQ4) To answer this question, before training the probes, we project the probing datasets $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ onto their $k$ largest principal components using PCA to obtain probes with substantially fewer parameters.
-3.5 Contrastive Activation Addition
-Our final set of experiments builds upon the findings of (Zhu et al., 2024), who showed that employing trained probes with inference time intervention (Li et al., 2023c, ITI) could enhance LMs’ performance on ToM tasks.
-We take a step further and ask: Can we enhance LMs’ performance by manipulating their activations without the need for training dedicated probes?
-(RQ5) To find an answer we use contrastive activation addition (Rimsky et al., 2023, CAA), an extension of activation addition (Turner et al., 2023, AA) that computes steering vectors to control LMs’ behaviour.
+Figure 3: Belief probing accuracy show similar patterns across all models: oracle belief representations generally form already in the first layers, while protagonist belief representations emerge at the intermediate layers.
+Moreover, probing accuracy increases with model size and, more crucially for smaller models, with fine-tuning.
+3.5 Activation Editing
+Prior work found that it is possible to manipulate models’ representations of beliefs by using (Li et al., 2023b, ITI), and that such interventions can improve LMs’ performance on ToM tasks.
+We take this further by asking whether a general “belief vector” can be distilled and injected into the models’ activations to strengthen their ToM abilities.
+To this end, we use contrastive activation addition (Rimsky et al., 2023, CAA), an extension of activation addition (Turner et al., 2023, AA) that computes steering vectors to control LMs’ behaviour.
 Steering vectors are computed as the average difference in residual stream activations between pairs of positive and negative instances of a specific behaviour.
 Formally, given a dataset $\mathcal{D}$ of triplets ($p$, $c_{p}$, $c_{n}$), where $p$ is a prompt, $c_{p}$ is a positive completion, and $c_{n}$ is a negative completion, CAA computes a mean difference vector $v^{md}_{l}$ for layer $l$ as:
-$v^{md}_{l}=\frac{1}{\absolutevalue{\mathcal{D}}}\sum_{p,c_{p},c_{n}}a_{l}(p,c_{p})-a_{l}(p,c_{n})$
-During inference, these steering vectors are multiplied with an appropriate coefficient $\alpha$ and added at every token position of the generated text after the user’s prompt.
-CAA has two main advantages over ITI: First, it eliminates the need to train probes.
+$v^{md}_{l}=\frac{1}{\absolutevalue{\mathcal{D}}}\sum_{p,c_{p},c_{n}\in\mathcal{D}}a_{l}(p,c_{p})-a_{l}(p,c_{n})$ (1)
+For example, in Figure 2, $p$ is the Story, $c_{p}$ could be the true belief, and $c_{n}$ the false belief.
+During inference, these steering vectors are multiplied by an appropriate coefficient $\alpha$ and added at every token position of the generated text after the prompt.
+CAA has two main advantages over ITI: First, it eliminates the need to train probes, making it computationally cheap.
+For example, for Llama2 70B, ITI needs to train 5,120 probes while CAA only needs to compute 80 vectors.
 Second, it operates at the residual stream level, making it easier to use than methods that intervene on specific attention heads like ITI.
 While CAA has been used to control alignment-relevant behaviour, such as hallucinations, refusal, and sycophancy (Rimsky et al., 2023), we are the first to apply it to enhance LMs’ ToM reasoning.
-This can be understood as isolating the direction in the LMs’ latent space corresponding to taking the perspective of another agent.
+The “belief vectors” (i.e.
+steering vectors) we obtain can be understood as isolating the direction in the LMs’ latent space corresponding to taking the perspective of another agent.
 To evaluate both base and fine-tuned LMs, we rank their answers to the ToM questions according to $p_{LM}(a|q)$ (Petroni et al., 2019).
-We adopt the Forward Belief task split used in (Zhu et al., 2024) to compute the steering vectors.
+For a fair comparison, we adopt the train/test Forward Belief split used in (Zhu et al., 2024) to compute and evaluate the steering vectors.
 Additionally, we evaluate the transferability of the CAA steering vectors by applying them to two other BigToM tasks: Forward Action and Backward Belief.
-We provide details about these tasks in Section A.1.1, and a more detailed explanation of how ITI works in Section A.5.
+We provide details about these tasks in Section A.1.1, and a more detailed explanation ITI in Section A.5.
 4 Results
-4.1 Effect of Model Size and Fine-tuning
+Figure 4: We compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $k=\{2,10,100,1000\}$ principal components.
+Results are for protagonist beliefs (for oracle see Figure 12).
+In general, it is possible to recover most of the original accuracy by training probes on a smaller number $k$ of principal components of the activations.
+Figure 5: Sensitivity of protagonist belief probing accuracy to different prompt variations. Results for Pythia are shown in Figure 9. Representations are brittle to prompt variations.
+Effect of Model Size and Fine-tuning
 Results from our study on model size and fine-tuning are shown in Figure 3.
-When considering oracle beliefs, probing accuracy rapidly converges to $100$, with larger models showing faster convergence rates.
-The smallest Pythia-70m that performs slightly worse but still achieves $95\%$ accuracy despite having less than $0.6\%$ of the parameters of Pythia-12B.
-This finding suggests that even small LMs can effectively represent beliefs from an omniscient perspective.
-For protagonist beliefs, accuracy also increases with model size, although there is a performance gap between Llama-2 and Pythia.
-For example, Llama2-13B reaches around $80\%$, while Pythia-12B achieves approximately $60\%$.
-This gap is likely due to Llama-2 being trained on nearly seven times more tokens than Pythia.
-The figure also shows that accuracy at early layers is particularly low across all models.
-We speculate that this is due to the initial coding strategy of LMs that uses the first layers to combine individual tokens into more semantically meaningful representations (Gurnee et al., 2023).
-Probes on fine-tuned LMs show significantly better accuracy with improvements of up to $29\%$ for Llama2-7B-chat and $26\%$ for Pythia-6.9B-chat with respect to their base version.
-Fine-tuned 7B LMs outperform (Llama-2) or are on par (Pythia) with twice as large base models (12/13B), highlighting the importance of fine-tuning in developing representations of others’ beliefs.
-This resonates with cognitive psychology findings that ToM development is closely linked to social communication (Tomasello, 2010; Sidera et al., 2018; Ma et al., 2023), which instruction-tuning and RLHF may help induce in language models.
-For larger LMs, the improvements from fine-tuning decrease as model size increases (Figure 6(a)).
-We characterise the relationship between probe accuracy and model size in Figure 6, where we consider the best probe accuracy for every LM, i.e.
-the highest accuracy among probes $\{g_{l}\}$ trained on $\{a_{l}\}$ for a LM $f$.
-For Llama-2 base, the best probe accuracy scales logarithmically with model size ($R^{2}=0.98$, Figure 6(b)), whereas for fine-tuned it scales linearly ($R=1.0$, cf.
-Figure 6(c)).
-For Pythia base, the best probe accuracy also scales logarithmically with model size ($R^{2}=0.96$, Figure 6(d)).
-4.2 Sensitivity to prompting
-Figure 3 compares protagonist probe accuracy across various prompt variations for different models, considering their architecture, size, and fine-tuning.
-As can be seen from the figure, providing the protagonist’s Initial Belief in the story yields higher probe accuracy compared to the Original prompt (Figure 1).
+For oracle beliefs, probing accuracy rapidly converges to $100$, with larger models showing faster convergence.
+Even the smallest Pythia-70m achieves $95\%$ accuracy.
+For protagonist beliefs, we notice a similar pattern across most models, where accuracy at early layers is particularly low and then increases at the intermediate layers.
+What happens at early layers is overfitting, which may be caused by spurious features introduced by the initial coding strategy of language models, where individual token representations are mixed together (Gurnee et al., 2023).
+We further discuss this in Section A.2.1.
+In general, probing accuracy increases with model size, although there is a performance gap between Llama-2 and Pythia.
+For example, Llama2-13B reaches around $80\%$ accuracy, while Pythia-12B achieves approximately $60\%$.
+This gap is likely due to Llama-2 being trained on nearly seven times more tokens than Pythia (cf.
+Table 2).
+Probes from fine-tuned LMs show significantly better accuracy, with improvements of up to $+29\%$ for Llama2-7B-chat (SFT + RLHF) and $+26\%$ for Pythia-6.9B-chat (SFT) compared to probes from their base version.
+The same probes outperform (Llama-2) or are on par (Pythia) with probes trained on twice as large base models (12/13B).
+This highlights a key role of fine-tuning in shaping belief representations in smaller LMs.
+The performance gap closes for the largest Llama2-70B, for which the improvements from fine-tuning are marginal.
+We characterise the relationship between probe accuracy and model size in Figure 7, using the best accuracy for each LM – i.e., the highest accuracy among probes ${g_{l}}$ trained on activations ${a_{l}}$ for model $f$.
+For Llama-2 base and Pythia base, probing accuracy scales logarithmically with model size (Figure 7(b), 7(d)), while for fine-tuned Llama-2 models, it scales linearly (Figure 7(c)).
+Control Tasks
+Figure 8(a) shows that probes trained on the control task consistently perform at random chance, confirming that higher probing accuracy in larger models meaningfully reflects a greater ability to extract ToM representations, rather than simply being a by-product of spurious correlations.
+For Llama models, the probes generally exhibit selectivity: they achieve high accuracy when probing for beliefs but remain at chance level on control tasks.
+Pythia’s overall accuracy is too low to allow for selectivity.
+Figure 4 shows probing accuracy on protagonist when training the probes on the top $k$ principal components of Llama-2’s internal activations.
+We provide results for Pythia in Figure 11, and for all models on oracle settings in Figure 12.
+We consider $k=\{2,10,100,1000\}$, spanning several orders of magnitude.
+[^3] For models with hidden dimensions smaller than $1000$, we skip this value.
+Results show that it is generally possible to recover most of the original accuracy by training probes on a smaller number $k$ of principal components of the activations.
+We also performed the first control experiment, this time only using the first $k=\{100,1000\}$ principal components.
+Figure 8(b) and 8(c) again show that probes trained on the control task consistently perform at random chance, confirming that probes are not fitting spurious patterns.
+Additionally, this suggests that belief representations are embedded in a low-dimensional subspace $\mathcal{B}$ spanned by the top $k$ eigenvectors $\{v_{1},\dots,v_{k}\}$ of the covariance matrix $\mathsf{C}=\mathbb{E}[(a-\mathbb{E}[a])(a-\mathbb{E}[a])^{\top}]$.
+Sensitivity to Prompting
+Figure 5 compares protagonist probe accuracy across various prompt variations for Llama-2 models.
+As can be seen from the figure, providing the protagonist’s Initial Belief in the story yields higher probe accuracy compared to the Original prompt.
 Accuracy for all the other prompt variations is generally lower than Original.
-On one hand, misleading prompts hurt performance across all models.
-This finding resonates with Webson & Pavlick (2022) who found that instruction-tuned models, despite being more robust, are still sensitive to misleading prompts.
+Misleading prompts hurt performance across all models.
+This finding resonates with Webson and Pavlick (2022), who found that instruction-tuned models, despite being more robust, are still sensitive to misleading prompts.
 On the other hand, Time Specification unexpectedly does not help in disambiguating belief states in different time frames, as we hypothesised in §3.4.
 Additionally, models show sensitivity to Random tokens placed before the belief statement.
+Pythia models show similar patterns, shown in Figure 9.
 Results for oracle beliefs are reported in Figure 10 and indicate that models maintain high accuracy.
 Misleading prompts slightly reduce performance to around $95\%$.
-In summary, these experiments show that LMs possess robust belief representations when taking an omniscient perspective, whereas their representations of others’ beliefs are more susceptible to prompt variations.
-4.3 Memorisation Effects in the Probes
-Figure 4 and Figure 8 show probe accuracies obtained by training a probe on the top $k$ principal components of the intermediate representations for protagonist and oracle, respectively.
-Specifically, we consider $k=\{2,10,100,1000\}$, spanning several orders of magnitude.
-For models with hidden dimensions smaller than $1000$, we skip this value.
-For all models, it is generally possible to recover most of the original accuracy by training probes on a number $k$ of principal components of the activations that is more than one order of magnitude smaller, indicating no strong evidence of memorisation in the probes.
-4.4 Contrastive Activation Addition
-We finally compare models’ accuracy on three BigToM tasks in Table 1.
+In summary, these experiments show that LMs possess robust belief representations when taking an omniscient perspective, whereas their representations of others’ beliefs are brittle to prompt variations.
+Model Method Forward Belief Forward Action Backward Belief
+TB FB Both TB FB Both TB FB Both
+Llama-2-7b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
+ITI $44_{+0}$ $44_{+0}$ $44_{+0}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$
+CAA $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}^{*}$ $71_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+27}}^{*}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}^{*}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}^{*}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}^{*}$ $74_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+30}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$
+Llama-2-7b-chat No int. $56$ $56$ $55$ $69$ $55$ $37$ $56$ $56$ $55$
+ITI $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $69_{+0}$ $55_{+0}$ $37_{+0}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+3}}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$
+CAA $70_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+14}}$ $72_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}^{*}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $69_{+0}$ $67_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $53_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}$ $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $84_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+27}}^{*}$ $57_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}^{*}$
+Llama-2-13b No int. $52$ $44$ $35$ $59$ $50$ $37$ $46$ $49$ $33$
+ITI $52_{+0}$ $45_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}$ $35_{+0}$ $64_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+5}}$ $61_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+11}}$ $46_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$ $48_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $42_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$
+CAA $85_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+33}}^{*}$ $88_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+44}}^{*}$ $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+31}}^{*}$ $71_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}^{*}$ $69_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+19}}^{*}$ $55_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+18}}^{*}$ $75_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+29}}^{*}$ $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+43}}^{*}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+26}}^{*}$
+Llama-2-13b-chat No int. $84$ $56$ $47$ $78$ $51$ $38$ $72$ $48$ $31$
+ITI $84_{+0}$ $65_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $78_{+0}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+7}}$ $47_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}^{*}$ $72_{+0}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $48_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+17}}$
+CAA $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}^{*}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+38}}^{*}$ $91_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+44}}^{*}$ $80_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}^{*}$ $71_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+20}}^{*}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}^{*}$ $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+25}}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+46}}^{*}$ $87_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+56}}^{*}$
+Llama-2-70b No int. $90$ $87$ $78$ $93$ $52$ $48$ $73$ $53$ $32$
+ITI $90_{+0}$ $90_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+3}}$ $78_{+0}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}$ $55_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+3}}$ $50_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $77_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+4}}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+5}}$ $37_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+5}}$
+CAA $99_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}^{*}$ $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}^{*}$ $95_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+17}}^{*}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}^{*}$ $80_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+28}}^{*}$ $73_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+25}}^{*}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+21}}$ $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+39}}^{*}$ $83_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+51}}^{*}$
+Llama-2-70b-chat No int. $69$ $75$ $56$ $86$ $56$ $52$ $63$ $59$ $52$
+ITI $69_{+0}$ $76_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $86_{+0}$ $56_{+0}$ $52_{+0}$ $63_{+0}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$
+CAA $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+23}}^{*}$ $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}^{*}$ $89_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+32}}^{*}$ $87_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}^{*}$ $75_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+19}}^{*}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+8}}^{*}$ $88_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+25}}$ $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+33}}^{*}$ $80_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+28}}$
+Table 1: Comparison of the effects of ITI (Li et al., 2023b) and CAA (Rimsky et al., 2023) on three tasks from BigToM (Gandhi et al., 2023).
+TB denotes a true belief task, whereas FB denotes a false belief task.
+The numbers represent accuracy scores, with the difference in performance compared to no intervention (No int.) indicated as subscripts.
+The asterisk (*) denotes a statistically significant difference from No int.
+based on a t-test with $p<0.05$.
+Results for Pythia are shown in Table 3.
+CAA outperforms ITI on all tasks.
+Contrastive Activation Addition
+We compare models’ accuracy on three BigToM tasks in Table 1 (Llama) and Table 3 (Pythia).
 Each model has been evaluated three times: without any intervention, using ITI, and using CAA.
 Hyperparameter details can be found in Section A.6.
-Note that we use steering vectors computed using the Forward Belief task for all three tasks to test their generalisability (Zhu et al., 2024).
-As can be seen from the table, performance without intervention is generally lower across tasks and model sizes, with the larger Llama-2-70B and Llama-2-70B-chat models exhibiting higher accuracy.
-Performance for Pythia models of different sizes does not change much, with the fine-tuned Pythia-6.9B-chat showing better performance on single true belief (TB) and false belief (FB) tasks but not on their conjunction (Both).
+Note that we use steering vectors computed using the Forward Belief task for all three tasks to test their generalisability.
+Performance without intervention is generally lower across tasks and model sizes, with the larger Llama-2-70B and Llama-2-70B-chat models exhibiting higher accuracy.
+Performance for Pythia models of different sizes does not change much, with the fine-tuned Pythia-6.9B-chat often showing better performance on single true belief (TB) and false belief (FB) tasks but not on their conjunction (Both).
 ITI demonstrates modest improvements over no intervention for Llama-2 models.
@@
 The only exception is Pythia-6.9B-chat, for which ITI is not always beneficial.
-CAA consistently delivers the most substantial accuracy improvements across all models and tasks, up to $+56$ for Llama-2-13B-chat on the (Backward Belief), which Gandhi et al.
+CAA consistently delivers the most substantial accuracy improvements across all models and tasks, up to $+56$ for Llama-2-13B-chat on the Backward Belief task, which Gandhi et al.
 have identified as the hardest task.
@@
 Smaller models like Pythia-70M and Pythia-410M also show significant gains with CAA, though the absolute performance is still lower than Llama-2.
-Overall, our results indicate that it is possible to effectively enhance ToM reasoning in LMs without needing to train any probe, which yields even improved results.
-Furthermore, we show that CAA steering vectors generalise well, yielding substantial performance gains across all ToM tasks.
-5 Limitations and Future Work
+To further demonstrate CAA’s effectiveness, we applied it while evaluating models on a control task where the causal event in the story is replaced by a random one that does not change the environment (e.g., A musician starts playing music while Noor is making the latte).
+Table 4 shows improved results for all models, indicating that CAA improves performance on ToM tasks without compromising the models’ ability on control tasks.
+Overall, our results indicate that it is possible to further enhance ToM reasoning in LMs in a computationally cheap way, without needing to train any probe.
+Furthermore, we show that the CAA steering vectors are general, yielding substantial performance gains across all ToM tasks.
+5 Discussion and Conclusion
+In this work, we conducted extensive experiments across 12 LMs to examine their internal representation of beliefs of self (oracle) and others (protagonist).
+Our experiments show similar emergence patterns across all the models we evaluated (RQ1): oracle belief representations generally form in the first layers, while for protagonist they emerge at the intermediate layers.
+Moreover, probing accuracy increases with model size and, more crucially for smaller models, with fine-tuning (RQ1) (Figure 3).
+While larger models show higher probing accuracy, this could be due to their higher dimensionality – at the same time increasing the number of learning parameters in the probes and offering more spurious patterns to fit.
+To control for this, we ran two experiments: one using randomly permuted labels, and one projecting activations onto their top-$k$ principal components to reduce probe size.
+Results show that high-dimensional probes cannot learn random label mappings (Fig.
+8), and that reduced representations retain most of the original accuracy (Fig.
+4, 11, 12).
+Together, these findings suggest that probes capture structured belief representations rather than spurious correlations (RQ2).
+We then explore if these representations are robust to prompt variations.
+Our experiments demonstrate that LMs possess robust belief representations when taking an omniscient perspective (Fig.
+10), whereas their representations of others’ beliefs are more brittle (RQ3), with probing accuracy decreasing for semantically neutral prompts (Fig.
+5, 9).
+Our final set of experiments shows that belief representations can be strengthened using CAA (RQ4).
+CAA steers model activations in a generalisable way, significantly improving performance across multiple ToM tasks while being computationally cheaper than ITI (Table 1, 3).
+For instance, with Llama-2-70B, ITI requires training 5,120 probes (64 attention heads $\times$ 80 layers), whereas CAA only needs 80 vectors, one per layer.
+In summary, our key takeaway is that while models can robustly represent beliefs from an omniscient perspective,
+representations of others’ beliefs improve with model size and fine-tuning, are structured yet brittle – but also easily steerable.
+Together, our findings suggest several promising directions for future work.
+Better understanding the similar emergence pattern of belief representations across LMs can inform architecture design and training strategies.
+Especially for smaller models, future work could explore how different types of fine-tuning (e.g., human feedback vs.
+synthetic data) influence the emergence of internal belief representations.
+Demonstrating that these representations are structured rather than spurious validates the use of probing as a meaningful tool to study how LMs’ represent beliefs of self and others, and encourages internal model analysis as part of evaluation pipelines.
+However, the brittleness of belief representations to prompts – particularly when attributing beliefs to others – suggests that the perspective-taking machinery needed for robust ToM reasoning remains fragile, and highlights the need for robustness benchmarks and new approaches to improve generalisation.
+Finally, our success with CAA shows that belief representations can be strengthened in a generalisable and efficient way, opening up opportunities for real-time model steering in socially grounded tasks.
+While CAA offers a post-hoc remedy, future research should also explore methods for directly embedding perspective-taking circuits into model architectures.
+Limitations
 Our study focused on expanding experiments from the model perspective, examining architectures, sizes, fine-tuning, and prompt design, all within the same dataset.
@@
 Nevertheless, our approach can be adopted to support new benchmarks or to evaluate newly released models as they become available.
-6 Conclusion
-Our study addresses a significant gap in understanding LMs by investigating their internal representation of mental states.
-We conducted an extensive benchmark involving various LM types, sizes, fine-tuning approaches, and prompt designs to examine the robustness of these representations.
-Our findings reveal that scaling LMs’ size and, in particular for smaller LMs, fine-tuning are key to developing representations of others’ beliefs.
-We are the first to demonstrate that such prompt variations influence model representations, and we also demonstrate the feasibility of enhancing models’ ToM reasoning by steering their activations without training any probe.
-Overall, our work contributes valuable insights into the factors influencing LMs’ mental state representations, shedding light on avenues for improving their performance in ToM tasks.
-Societal Impact
-While our work is foundational and remains distant from specific applications with direct societal impact, it’s important to recognise the ethical implications of modelling and predicting mental states.
-Handling sensitive aspects of individuals’ inner experiences and emotions requires careful consideration to avoid reinforcing biases or misunderstanding psychological nuances.
-Author Contributions Statement
-M. Bortoletto designed and directed the project, performed the experiments, analysed the results, and wrote the manuscript.
-C. Ruhdorfer provided feedback on the experimental results and contributed to editing the final manuscript.
-L. Shi contributed to proofreading the final manuscript.
-A. Bulling supervised the project and contributed to the writing of the manuscript.
+Finally, while in this work we focused on beliefs, our experimental approach can be adapted to investigate how LMs represent desires, emotions, intentions, or preferences.
+Future research exploring other types of mental states can use our findings to determine whether similar or distinct patterns emerge.
 Acknowledgements
@@
 Ruhdorfer.
-The authors would like to especially thank Hsiu-Yu Yang and Anna Penzkofer for their support and the numerous insightful discussions.
-We used PyTorch (Paszke et al., 2019) to manipulate tensors, Transformers (Wolf et al., 2020) to evaluate the models and NNsight (Fiotto-Kaufman, ) to perform activation interventions.
-We performed our data analysis using NumPy (Harris et al., 2020), Pandas (pandas development team, 2020; Wes McKinney, 2010), and SciPy (Virtanen et al., 2020).
-We made our plots using Matplotlib (Hunter, 2007) and seaborn (Waskom, 2021).
 Appendix A Appendix
@@
 (2024) and Gandhi et al.
-(2023), we focused on the 6 most important conditions, corresponding to true and false beliefs on the following three tasks:
+(2023) we focused on the 6 most important conditions, corresponding to true and false beliefs on the following three tasks:
   • Forward Belief: given the protagonist’s percepts of the causal event, infer their belief: $P(\mathrm{belief}|\mathrm{percept})$.
@@
 The dataset was released under the MIT license and can be accessed at https://github.com/cicl-stanford/procedural-evals-tom.
-We report one example for each task in the boxes below, where the text defining true belief or false belief task is shown in blue and red, respectively.
-Example of Forward Belief Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor sees her coworker swapping the milk.
-Noor does not see her coworker swapping the milk.
-Does Noor believe the milk pitcher contains oat milk or almond milk?
-a) Noor believes the milk pitcher contains oat milk.
-b) Noor believes the milk pitcher contains almond milk.
-Example of Forward Action Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor sees her coworker swapping the milk.
-Noor does not see her coworker swapping the milk.
-What will Noor do?
-a) Noor will make the cappuccino using the milk in the pitcher.
-b) Noor will open the fridge once again to take out the oat milk and replace the almond milk with oat milk.
-Example of Backward Belief Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor opens the fridge again and reaches for the oat milk.
-Noor makes the cappuccino using the milk in the pitcher.
-Does Noor believe the milk pitcher contains oat milk or almond milk?
-a) Noor believes the milk pitcher contains oat milk.
-b) Noor believes the milk pitcher contains almond milk.
+We report one example for each task in Example A.9, A.9, and A.9, where the text defining true belief or false belief task is shown in blue and red, respectively.
 A.1.2 Linear probes
 Our probing approach is illustrated in Figure 6.
-We cache activations both at the attention and residual stream level.
-In our experiments, we use residual stream activations.
-To perform ITI and compare it to CAA, we also cache attention heads’ activations.
-We trained the probes using the L-BFGS solver (Liu & Nocedal, 1989) with L2 penalty with inverse of regularisation strength $10$ for a maximum of $1000$ iterations.
+For our experiments, we cache activations at the residual stream level.
+To perform ITI and compare it to CAA, we also cache attention heads activations.
+We trained the probes using the L-BFGS solver (Liu and Nocedal, 1989) with L2 penalty with inverse of regularisation strength $10$ for a maximum of $1000$ iterations.
 We use zero as random seed.
@@
 For all the models, we set the temperature to zero.
-Table 2: The 12 models used in this work. The checkmark indicates we additionally study the fine-tuned (Chat) version of the model.
-LM Size + Chat Tokens $d_{model}$ Layers
-Llama-2 (Touvron et al., 2023) 7B 2T 4096 32
+LM Size + SFT + RLHF Tokens $d_{model}$ Layers
+Llama-2 7B 2T 4096 32
 13B 2T 5120 40
 70B 2T 8192 80
-Pythia (Biderman et al., 2023) 70M 300B 512 6
+Llama-2-chat 7B 2T 4096 32
+13B 2T 5120 40
+70B 2T 8192 80
+Pythia 70M 300B 512 6
 410M 300B 1024 24
@@
 6.9B 300B 4096 32
-12B 300B 5120 40
+12B 300B 5120 36
+6.9B 300B 4096 32
+Table 2: The 12 models used in this work.
 A.1.4 Examples of prompt variations
-Default prompt Story: Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor does not see her coworker swapping the milk.
-Belief: Noor believes the milk pitcher contains oat milk.
-Random prompt Story: Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor does not see her coworker swapping the milk.
-Belief: 장士 decided [_ countries sections behoSOUR gminy bef Noor believes the milk pitcher contains oat milk.
-Misleading prompt Story: Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor does not see her coworker swapping the milk.
-Belief: Noor believes the milk pitcher contains oat milk.
-Belief: The water valve is closed.
-Time specification prompt Story: Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor does not see her coworker swapping the milk.
-Belief: In the end, Noor believes the milk pitcher contains oat milk.
-Initial belief prompt Story: Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-Noor believes that the milk pitcher contains oatmilk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor does not see her coworker swapping the milk.
-Belief: Noor believes the milk pitcher contains oat milk.
+Example A.9 shows an example of Original prompt.
+Examples of prompt variations are provided in Example A.9 (Random), Example A.9 (Misleading), Example A.9 (Time Specification), and Example A.9 (Initial Belief).
 A.2 Model size and fine-tuning
@@
 For Pythia base, the best probe accuracy also scales logarithmically with model size ($R^{2}=0.96$, Figure 7(d)).
+A.2.1 Overfitting Issues
+Figure 3 also that probing accuracy at early layers is particularly low across all models, performing even worse than random.
+This happens due to overfitting, which may be caused by spurious features introduced by the initial coding strategy of language models, where individual token representations are mixed together (Gurnee et al., 2023).
+We also identified the same issue when reproducing the results in Zhu et al.
+(2024), who address it by manually clipping all accuracies below random chance to 50%.
+[^4] https://github.com/Walter0807/RepBelief/blob/0fc86396f2f0a998643ea01786eb3db4dd20ff9c/probe.py#L60 Since probing experiments require training a large number of probes for each model, both we and Zhu et al.
+(2024) trained each probe for the same fixed number of epochs (1,000).
+However, for activations from the earlier layers, overfitting occurs very quickly - often within the first 10 iterations.
+We ran an experiment with Llama2-7B-chat, reducing training to fewer than 10 iterations, and found that the probes performed at random chance.
+Therefore, to fully resolve this issue, we would need to choose the number of training epochs for each probe individually.
+This would likely flatten the observed "U" shape in the results.
+However, this process would be computationally expensive and does not contribute to our main research questions.
+Rather than artificially adjusting accuracies to 50%, we prefer to present the results as they are.
 Figure 7: To characterise the relationship between probe accuracy and model size we consider the best probe accuracy for every LM, i.e.
@@
 (d) Logarithmic fit for Pythia base.
+(a) Original activations.
+(b) $k=1000$ largest principal components of the activations.
+(c) $k=100$ largest principal components of the activations.
+Figure 8: Comparison between accuracy on belief probing and accuracy obtained on a control task.
 A.3 Sensitivity to prompting
+Accuracy on protagonist belief probing for Pythia models is shown in Figure 9.
 Accuracy on oracle belief probing for different prompt variations are reported in Figure 10.
 Figure 9: Sensitivity of protagonist belief probing accuracy to different prompt variations.
+Figure 10: Sensitivity of protagonist belief probing accuracy to different prompt variations.
 A.4 Dimensionality reduction
+Probing accuracy obtained by Pythia models for the protagonist setting is reported in Figure 11.
 Oracle probe accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components are shown in Figure 12.
+Figure 11: We compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components.
+For Pythia: All(70m) = $512$, All(410m) = $1024$, All(1b) = $2048$, All(6.9b) = $4096$, All(12b) = $5120$.
+Results for oracle are shown in Figure 12.
 Figure 12: (Oracle) To investigate potential memorisation in the probes, we compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components.
@@
 A.5 Inference-time intervention
-Inference-time intervention (Li et al., 2023c, ITI) employs a two-step process.
+Inference-time intervention (Li et al., 2023b, ITI) employs a two-step process.
 First, it trains a probe for each attention head across all layers of a LM.
@@
 Formally, ITI can be defined as an additional term to the multi-head attention:
-$x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\left(\mathrm{Att}_{l}^{h}(P_{l}^{h}x_{l}){\color[rgb]{0.5,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,1}\pgfsys@color@cmyk@stroke{0.50}{1}{0}{0}\pgfsys@color@cmyk@fill{0.50}{1}{0}{0}\ +\ \alpha\sigma_{l}^{h}\theta_{l}^{h}}\right)$
+$x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\left(\mathrm{Att}_{l}^{h}(P_{l}^{h}x_{l}){\color[rgb]{0.56,0.27,0.52}\definecolor[named]{pgfstrokecolor}{rgb}{0.56,0.27,0.52}\ +\ \alpha\sigma_{l}^{h}\theta_{l}^{h}}\right)$
 where $x_{l}$ is the residual stream at layer $l$, $H$ is the number of attention heads, $\alpha\in\mathbb{R}^{+}$ is a coefficient, $\sigma_{l}^{h}$ is the standard deviation of activations along the direction identified by the probe trained on attention head $h$ at layer $l$, and $\theta_{l}^{h}$ is zero ofr not-selected attention heads.
-A.6 Activation editing hyperparameters
-Table 3 reports results obtained on the three BigToM tasks with the hyperparameters used for ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023).
-We report an example of prompt used for evaluation in the box below.
-Prompt used for downstream evaluation (Section 4.4) Answer the questions based on the context.
-Keep your answer concise, few words are enough, maximum one sentence.
-Answer as ’Answer:<option>)<answer>’.
-Story: Noor is working as a barista at a busy coffee shop.
-Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
-Noor grabs a milk pitcher and fills it with oat milk.
-A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
-Noor does not see her coworker swapping the milk.
-Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
-Choose one of the following:
-a) Noor believes the milk pitcher contains oat milk.
-b) Noor believes the milk pitcher contains almond milk.
-Answer:
-Table 3: Comparison between ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023) on the three BigToM tasks (Gandhi et al., 2023).
-For ITI, the subscript indicates the value of the coefficient used.
-For CAA, the subscript indicates first the value of the coefficient used and second the layer at which intervention takes place.
+A.6 Activation editing
+Table 3 reports results obtained on the three BigToM tasks with the corresponding hyperparameters used for ITI (Li et al., 2023b) and CAA (Rimsky et al., 2023).
+We report an example of prompt used for evaluation in Example A.9.
+Table 4 shows the accuracy obtained by using CAA on the Forward Belief True Control task in BigToM.
+On this control task, CAA produced improved results for all model, proving that CAA not only improves performance on ToM tasks, but also does not degrades the models’ ability to perform other tasks.
 Model Method Forward Belief Forward Action Backward Belief
-TB FB TB $\land$ FB TB FB TB $\land$ FB TB FB TB $\land$ FB
+TB FB Both TB FB Both TB FB Both
 Llama-2-7b No int. $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$ $44$
@@
 CAA $54_{2.0,0}$ $64_{2.0,9}$ $54_{2.0,0}$ $60_{2.0,11}$ $58_{2.0,11}$ $55_{2.0,12}$ $54_{2.0,0}$ $67_{2.0,10}$ $54_{2.0,0}$
+Table 3: Activation intervention: comparison between ITI (Li et al., 2023b) and CAA (Rimsky et al., 2023).
+For ITI, the subscript indicates the value of the coefficient $\alpha_{\mathrm{ITI}}$ used: $\mathrm{Acc}_{\alpha_{\mathrm{ITI}}}$.
+For CAA, the subscript indicates first the value of the coefficient $\alpha$ used and second the layer $l$ at which intervention takes place: $\mathrm{Acc}_{\alpha_{\mathrm{CAA}},l}$.
+Model Method Control CAA Parameters
+Llama-2-7b No int. $44$
+CAA $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}$ $2.0$, $11$
+Llama-2-7b-chat No int. $56$
+CAA $70_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+14}}$ $1.0$, $11$
+Llama-2-13b No int. $52$
+CAA $85_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+33}}$ $2.0$, $12$
+Llama-2-13b-chat No int. $84$
+CAA $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}$ $1.0$, $12$
+Llama-2-70b No int. $90$
+CAA $99_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$ $2.0$, $16$
+Llama-2-70b-chat No int. $69$
+CAA $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+23}}$ $1.5$, $18$
+Pythia-70m No int. $41$
+CAA $62_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+21}}$ $1.0$, $2$
+Pythia-410m No int. $48$
+CAA $67_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+19}}$ $2.0$, $4$
+Pythia-1b No int. $44$
+CAA $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+15}}$ $2.0$, $8$
+Pythia-6.9b No int. $44$
+CAA $56_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $1.5$, $12$
+Pythia-6.9b-chat No int. $55$
+CAA $68_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}$ $1.5$, $15$
+Pythia-12b No int. $44$
+CAA $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $2.0$, 00
+Table 4: Results for CAA Rimsky et al. (2023) on the Forward Belief True Control task in BigToM Gandhi et al. (2023). Numbers indicate accuracy scores, with differences (CAA -- No int.) as subscripts.
 A.7 Compute resources
@@
 A.8 Code
-Our code will be made public under the MIT licence at https://git.hcics.simtech.uni-stuttgart.de/public-projects/mental-states-in-LMs.
+Our code is provided as supplementary material and it will be made public under the MIT licence at www.this-is-a-placeholder.com.
+A.9 Societal impact
+While our work is foundational and remains distant from specific applications with direct societal impact, it’s important to recognise the ethical implications of predicting and editing mental state representations.
+Handling sensitive aspects of individuals’ inner experiences and emotions requires careful consideration to avoid reinforcing biases or misunderstanding psychological nuances.
+As LMs begin to encode aspects of ToM, there’s a risk that over-interpreting these capabilities could lead to misplaced trust – especially in real-world applications requiring nuanced social reasoning, such as education, healthcare, or mental health support.
+Furthermore, while techniques like CAA show promise for steering internal representations, they also potentially introduce new ethical challenges.
+Manipulating a model’s internal states, especially in ways that affect social reasoning, requires transparency and caution to avoid unintended consequences such as bias amplification or fairness issues.
+Future work should consider not only improving technical performance but also developing safeguards and evaluation frameworks to ensure responsible use of ToM-like abilities in LMs.
+Example 1: Forward Belief Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor sees her coworker swapping the milk.
+Noor does not see her coworker swapping the milk.
+Does Noor believe the milk pitcher contains oat milk or almond milk?
+a) Noor believes the milk pitcher contains oat milk.
+b) Noor believes the milk pitcher contains almond milk.
+Example 2: Forward Action Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor sees her coworker swapping the milk.
+Noor does not see her coworker swapping the milk.
+What will Noor do?
+a) Noor will make the cappuccino using the milk in the pitcher.
+b) Noor will open the fridge once again to take out the oat milk and replace the almond milk with oat milk.
+Example 3: Backward Belief Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor opens the fridge again and reaches for the oat milk.
+Noor makes the cappuccino using the milk in the pitcher.
+Does Noor believe the milk pitcher contains oat milk or almond milk?
+a) Noor believes the milk pitcher contains oat milk.
+b) Noor believes the milk pitcher contains almond milk.
+Example 4: Default prompt Story: Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor does not see her coworker swapping the milk.
+Belief: Noor believes the milk pitcher contains oat milk.
+Example 5: Prompt variation – Random Story: Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor does not see her coworker swapping the milk.
+Belief: 장士 decided [_ countries sections behoSOUR gminy bef Noor believes the milk pitcher contains oat milk.
+Example 6: Prompt variation – Misleading Story: Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor does not see her coworker swapping the milk.
+Belief: Noor believes the milk pitcher contains oat milk.
+Belief: The water valve is closed.
+Example 7: Prompt variation – Time specification Story: Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor does not see her coworker swapping the milk.
+Belief: In the end, Noor believes the milk pitcher contains oat milk.
+Example 8: Prompt variation – Initial belief Story: Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+Noor believes that the milk pitcher contains oatmilk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor does not see her coworker swapping the milk.
+Belief: Noor believes the milk pitcher contains oat milk.
+Example 9: Prompt used for downstream evaluation Answer the questions based on the context. Keep your answer concise, few words are enough, maximum one sentence. Answer as ’Answer:<option>)<answer>’.
+Story: Noor is working as a barista at a busy coffee shop.
+Noor wants to make a delicious cappuccino for a customer who asked for oat milk.
+Noor grabs a milk pitcher and fills it with oat milk.
+A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task.
+Noor does not see her coworker swapping the milk.
+Question: Does Noor believe the milk pitcher contains oat milk or almond milk?
+Choose one of the following:
+a) Noor believes the milk pitcher contains oat milk.
+b) Noor believes the milk pitcher contains almond milk.
+Answer: